### v2

# Spectral theory of isogeny graphs

Abstract
We consider finite graphs whose vertexes are supersingular elliptic curves, possibly with level structure, and edges are isogenies. They can be applied to the study of modular forms and to isogeny based cryptography. The main result of this paper says that these graphs have the Ramanujan property, which means that the eigenvalues of their adjacency matrices are as small as possible. We also study the asymptotic distribution of the eigenvalues of the adjacency matrices, the number of connected components, the automorphisms of the graphs, and the connection between the graphs and modular forms.
1 Introduction
Given two distinct prime numbers $p$ and $\ell$, supersingular isogeny graphs are finite graphs whose vertexes are isomorphism classes of supersingular elliptic curves defined over a field of characteristic $p$, possibly enriched with some level structure, and edges are degree $\ell$ isogeny, see Definitions 1.1 and 1.2. The number of vertexes of these graphs grows linearly in $p$.
Theorems 1.4 and 1.6, our main results, give information about the spectrum of the adjacency matrices of these graphs. They rely on algebraic geometry constructions.
The spectrum of the adjacency matrix is not a complete invariant of a graph, indeed there are graphs, sometimes called cospectral mates, which are non-isomorphic but still their adjacency matrices have the same spectrum. However, results from spectral graph theory gather important information about the geometry of the graph only out of the spectrum of the adjacency matrix. This is why our work provides a better understanding of isogeny graphs.
Isogeny graphs were first studied by Mestre [37] in the 80’s. His goal was to study modular forms, in particular to compute eigenforms out of eigenvectors of adjacency matrices of isogeny graphs. This approach has been recently made very practical in [14]. Our Theorems 6.5.2, 6.5.5 generalize [37, Theorem 2.1], and we hope they lead to possible extensions of Mestre’s “Méthode des graphes”, even though an analogue for formula (1) in loc. cit. is needed.
In the 90’s people from graph theory were looking for explicit examples of graph with optimal spectral gap, and consequently optimal expansion constant and mixing time. Surprisingly, classical isogeny graphs, i.e. without level structure, provided such examples! These facts are discussed in Section 1.2, where we also show, as corollary of our main results, that also isogeny graphs with level structure have this property
More recently, isogeny graphs started to play an important role in cryptography, as many protocols from isogeny based cryptography rely on their features. For instance, in [7] information about the spectrum of isogeny graphs with Borel level structure is used to prove Statistical secure Zero Knowledge Proof. This is discussed in Section 1.3
1.1 Main Definitions and Results
Definition 1.1 (Level structure on elliptic curves)
Fix a positive integer $N$ and a subgroup $H$ of $\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})=\mathrm{Aut}((\mathbb{Z}/N\mathbb{Z})^{2})$. For each field $k$ whose characteristic does not divide $N$ and each elliptic curve $E/k$, a level $H$ structure on $E$ is an isomorphism $\phi\colon(\mathbb{Z}/N\mathbb{Z})^{2}\to E[N]$ considered up to composition with an element of $H$, i.e. we consider two isomorphism $\phi$ and ϕsuperscriptitalic-$\phi^{\prime}$ equivalent if there exists an element $h$ in $H$ such that $\phi=\phi^{\prime}\circ h$.
Sometimes level $H$ structures have a more explicit interpretations, as illustrated below.
  • Trivial level structure When $H=\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$, there is a unique level structure on every elliptic curve;
  • Borel level structure When $H=\{\left(\begin{smallmatrix}*&0\\ *&*\\ \end{smallmatrix}\right)\}$ is the subgroup of lower triangular matrices, an $H$ level structure is equivalent to the choice of cyclic a subgroup of order $N$ in $E[N]$;
  • Full level structure When $H=\{\mathrm{Id}\}$, an $H$ structure is equivalent to the choice of a basis of $E[N]$;
  • Split Cartan level structure When $H=\{(\begin{smallmatrix}*&0\\ 0&*\end{smallmatrix})\}$, a level structure is equivalent to the choice of a ordered pair of cyclic a subgroups $C_{1},C_{2}<E[N]$ having order $N$ and trivial intersection. This level structure gives a graph isomorphic to a graph with Borel level structure, see Section 2.4, so we will not discuss it in details. It is also possible to take the normalizer of the Cartan, this correspond to take a non-ordered pair of cyclic subgroup, the corresponding graph is a quotient of the graph with Cartan level structure.
  • Torsion point level structure When $H=\{(\begin{smallmatrix}*&0\\ *&1\end{smallmatrix})\}$, an $H$ level structure is equivalent to the choice of a point of order exactly $N$;
  • Non split Cartan level structure It is defined by (the unique up to conjugation) non-split Cartan subgroups of $\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$. Details are given [19] and in [42] these structures are interpreted as “necklaces” of subgroups of $E[N]$ for $N$ prime.
Fix $(E_{1},\phi_{1})$ and $(E_{2},\phi_{2})$, where $E_{1},E_{2}$ are elliptic curves over a common field $k$, and $\phi_{i}$ is a level $H$ structure on $E_{i}$. A morphism $\alpha\colon(E_{1},\phi_{1})\to(E_{2},\phi_{2})$ is an isogeny $\alpha\colon E_{1}\to E_{2}$ such that $\alpha\circ\phi_{1}=\phi_{2}$ as level $H$ structures on $E_{2}$, or equivalently such that there exists an element $h\in H$ satisfying $\alpha\circ\phi_{1}=\phi_{2}\circ h$. The degree of such a morphism is the degree of the corresponding isogeny. A morphism is an isomorphism if the corresponding isogeny is an isomorphism, i.e. it has degree one.
Definition 1.2 (Supersingular isogeny graph)
Fix a positive integer $N$, a subgroup $H$ of $\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$ and distinct prime numbers $p,\ell$ not dividing $N$.
The isogeny graph with level structure $G=G(p,\ell,H)$ is the graph with:
  • vertices $V=\{(E_{1},\phi_{1}),\dots,(E_{r},\phi_{r})\}$ a set of representatives of isomorphism classes of supersingular elliptic curves $E/\overline{\mathbb{F}_{p}}$ with a level $H$ structure $\phi$.
  • edges: given vertexes $(E_{i},\phi_{i})$ and $(E_{j},\phi_{j})$, edges between them are degree $\ell$ morphisms $(E_{i},\phi_{i})\to(E_{j},\phi_{j})$, modulo automorphisms of $(E_{j},\phi_{j})$.
We denote $A={(a_{ij})_{i,j}}$ the adjacency matrix of $G$, namely the matrix whose entries $a_{ij}$ are the number of edges $(E_{j},\phi_{j})\to(E_{i},\phi_{i})$.
Remark 1.3
Suppose $E/\overline{\mathbb{F}}_{p}$ is a supersingular elliptic curve with an automorphism $u$, and that $\phi\colon(\mathbb{Z}/N\mathbb{Z})^{2}\to E[N]$ is a level $H$ structure on $E$. Then, the pairs $(E,\phi)$ and $(E,u\circ\phi)$ are always isomorphic, hence there is one vertex $(E_{i},\phi_{i})$ of $G(p,\ell,H)$ representing both. Nevertheless $u$ does not always define an automorphism of $(E_{i},\phi_{i})$: it does if and only if the endomorphism $\phi^{-1}\circ u\circ\phi$ of $(\mathbb{Z}/N\mathbb{Z})^{2}$ lies in $H$. In particular, if $\left(\begin{smallmatrix}-1&\\ &-1\\ \end{smallmatrix}\right)\notin H$, then $-1$ is not an automorphism of $(E,\phi)$ even though $(E,\phi)\cong(E,-\phi)$.
In the context of the above definition, given a vertex $(E_{i},\phi_{i})$, taking the kernel of isogenies gives a bijection between cyclic subgroup of cardinality $\ell$ of $E_{i}[\ell]$, and edges coming out of the vertex $(E_{i},\phi_{i})$. In particular there are exactly $\ell{+}1$ edges coming out of each vertex.
The graph $G$ might not be connected. For every connected component $G_{i}$, consider the vector $v_{i}$ in $\mathbb{C}^{V}$ obtained as formal sum of the vertex of $G_{i}$. Then ${}^{t}Av_{i}=(\ell+1)v_{i}$, where ${}^{t}$ denotes the transpose. This shows that $\ell{+}1$ is an eigenvalue of $A$.
Our first main result is the following
Theorem 1.4
With the notation of Definition 1.2, if $H$ contains the scalar matrices and $\det(H)=(\mathbb{Z}/N\mathbb{Z})^{\times}$, then the graph $G(p,\ell,H)$ is connected, its adjacency matrix $A$ is diagonalizable, the eigenvectors are real, the eigenvalue $\ell+1$ has multiplicity one, and all the other eigenvalues are contained in the open Hasse interval $(-2\sqrt{\ell},2\sqrt{\ell})$.
The above result covers the case of graphs with Borel, Cartan (both split and non-split) and trivial level structure. We notice that the graph with trivial level structure coincides with the classical isogeny graphs.
When the graph contains pairs $(E,\phi)$ with non-trivial automorphisms (i.e. automorphisms not induced by ±$\pm 1\in\mathrm{Aut}(E)$), the adjacency matrix $A$ is not symmetric, hence the fact the spectrum is real requires some non-trivial argument.
When $\det(H)$ is strictly smaller than $(\mathbb{Z}/N\mathbb{Z})^{\times}$, we need to introduce some further notations to describe the connected components of the graphs, and their partitions. Let $\mu_{N}^{\times}(\overline{\mathbb{F}}_{p})$ be the set of primitive $N$-th root of unity in $\overline{\mathbb{F}}_{p}$. This is a principal homogeneous space for the right action of $(\mathbb{Z}/N\mathbb{Z})^{\times}$ given by $\zeta\cdot a=\zeta^{a}$. The group $\det(H)$ is a subgroup of $(\mathbb{Z}/N\mathbb{Z})^{\times}$, so it also acts on $\mu_{N}^{\times}(\overline{\mathbb{F}}_{p})$ and we can form the quotient $R_{H}:=\mu_{N}^{\times}(\overline{\mathbb{F}}_{p})/\det H$.
Definition 1.5 (Weil invariant of a level structure)
Consider an elliptic curve with $H$ level structure $(E,\phi)$. Let $w$ be the Weil pairing on the $N$-torsion of $E$ and let
$w(\phi)=w(\phi(\begin{smallmatrix}1\\ 0\end{smallmatrix}),\phi(\begin{smallmatrix}0\\ 1\end{smallmatrix}))\quad\in R_{H}\,.$
As $\phi$ is defined only modulo the action of $H$, the invariant $w(\phi)$ is an element of the quotient $R_{H}$. We call this invariant the Weil invariant of the level structure.
The Weil invariant gives an obstruction to $G=G(p,\ell,H)$ being connected: if two vertices $v_{1},v_{2}$ are connected by a degree $\ell$ isogeny, then [46, Chapter III, Proposition 8.2] implies that their corresponding Weil invariants are connected by the action of $\ell$, i.e. $w(v_{2})=w(v_{1})^{\ell}$. Hence in a connected component of $G$, the Weil invariant has image an orbit of the action of $\ell$ on $R_{H}$. Let $\{C_{1},\dots,C_{n}\}$ be the orbits of $\ell$ acting on $R_{H}$ and for each $i$ we denote
$G_{i}:=w^{-1}(C_{i})$
which is a subgraph of $G$ by the previous argument. Our second main result generalizes Theorem 1.4.
Theorem 1.6
With the notation of Definition 1.2, let $G=G(p,\ell,H)$, and let $G_{1},\ldots,G_{n}$ be the subgraphs of $G$ defined above.
Connected components
Each $G_{i}$ is connected, i.e. the graph $G$ has $n$ connected components. Let $N(H)$ be the normalizer of $H$ in $\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$. If $p$, $\ell$ and $\det(N(H))$ generate $(\mathbb{Z}/N\mathbb{Z})^{\times}$, then all $G_{i}$’s are isomorphic.
Spectrum of the adjacency matrix
Denote $k$ the order of $\ell$ in $(\mathbb{Z}/N\mathbb{Z})^{\times}/\det H$, and $k^{\prime}$ the smallest positive integer such that $\ell^{k^{\prime}}\mathrm{Id}\in H$. The adjacency matrix $A_{i}$ of $G_{i}$ is diagonalizable and, for each $k$-th root of unity $\zeta$, the number $(\ell{+}1)\zeta$ is an eigenvalue of $A_{i}$ of multiplicity one. The other eigenvalues of $A_{i}$ are complex numbers with angle in $\mathbb{Z}\tfrac{\pi}{k^{\prime}}$ and absolute value in the open Hasse interval $(-2\sqrt{\ell},2\sqrt{\ell})$.
Theorem 1.6 applies to the case of full level structure, where the adjacency matrix has non-real eigenvalues. In this case $N(H)=\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$, hence all connected components are isomorphic. We also have that $k=k^{\prime}$ is the multiplicative order of $\ell$ in $(\mathbb{Z}/N\mathbb{Z})^{\times}$, and the number of connected components is $n=\phi(N)/k$. Our description of the connected components also answers questions and conjectures from [18].
We can also apply Theorem 1.6 to the isogeny graphs with torsion point level structure, namely $H=\{\left(\begin{smallmatrix}*&0\\ *&1\\ \end{smallmatrix}\right)\}$. In this case $\det H=(\mathbb{Z}/N\mathbb{Z})^{\times}$, hence $G$ is connected and $k=1$. One might have $k^{\prime}>1$, and indeed Corollary 1.11 implies that for $p$ big enough the adjacency matrix has non-real eigenvalues.
Remark 1.7 (Multipartite graphs)
Given a finite connected directed graph $G=(V,E)$, a $k$-multipartition is a partition of $V$ into $k$ disjoint subsets $V_{j}$ such that vertexes of $V_{j}$ are connected only to vertexes of $V_{j+1}$. 2-partite graphs are called bipartite. When $G$ is $d$-regular, this is related to the spectrum of the adjacency matrix $A$ of $G$ in the following way. Let $u_{j}=\sum_{v\in V_{j}}v$, and $U$ the span of $\{u_{1},\dots,u_{k}\}$ in $\mathbb{C}^{V}$. Then $U$ is stabilized by $A^{t}$, and $A^{t}$ restricted to $U$ acts as $d$ times a cyclic permutation, hence the spectrum of $A$ contains $d$ times the group of $k$-th root of unity.
Then Weil invariant gives a $k$-multipartion of the vertexes of $G_{i}$, and by the above discussion this is a $k$-multipartion of $G_{i}$; the existence of this partition implies the existence of the eigenvalues $(\ell+1)\zeta$’s appearing in the statement of Theorem 1.6. Theorem 1.6 also says that there are no other eigenvalues of module $\ell+1$, hence this partition can not be refined.
Organization of the paper
In Section 2, we reduce the proof of Theorems 1.4 and 1.6 to Theorem 2.3.6.
Sections 3, 4 and 5 are devoted to set-up a more general framework to study isogeny graphs, and to prove a slightly weaker version of Theorem 2.3.6(= Theorem 5.7), namely that the moduli of some eigenvalues are contained in the closed Hasse interval $[-2\sqrt{\ell},2\sqrt{\ell}]$. These Sections rely on more advanced algebraic geometry notions.
In Section 6 we develop the connection between isogeny graphs and modular forms; this connection is used to streghtened the above mentioned result, namely to show the containment in the open Hasse interval $(-2\sqrt{\ell},2\sqrt{\ell})$, and to prove Corollary 1.11.
Trough the paper, we keep track of automorphisms of the graphs. We relate them to autormophisms of modular curves and modular forms, such as the Fricke involution and Atkin-Lehner automorphisms. These results are not used in the proof of our main theorems, but we think they can be useful for further developments.
1.2 Ramanujan graphs and expander sequences
In this section we discuss the implication of our results from the point view of graph theory. We refer the reader to the textbooks [15, 29], the papers [9, 25] and references therein for detailed discussions of the concepts introduced here.
Let $G$ be a $d$-regular non-bipartite (see Remark 1.7) connected finite graph with symmetric adjacency matrix $A$. The spectrum of $A$ contains the eigenvalue $d$, called trivial eigenvalue, with multiplicity one. All other eigenvalues are called non-trivial and are contained in the interval $(-d,d)$ ([15, Proposition 1.1.2]). The spectral gap is the minimum of $d-|\lambda|$, where $\lambda$ runs among all non-trivial eigenvalues. The spectral gap can be used, among the other things, to bound the diameter, the expansion constant and the mixing time of a graph, see [15, 29].
A graph is called Ramanujan if all non-trivial eigenvalues of $A$ are contained in the Hasse interval $[-2\sqrt{d-1},2\sqrt{d-1}]$. Alon-Boppana inequality implies that Ramanujan graphs have the largest possible spectral gap among big graphs ([29, Section 5.2], [15, Section 1.3], [9, Introduction]). A key result, conjectured by Alon and proven in [25] and [9], says the following: fixed a positive number $\varepsilon$, using the uniform distribution on the set of $d$-regular simple graphs with $n$ vertexes, the probability that all non trivial eigenvalues of the adjacency matrix lie in the interval $[-2\sqrt{d-1}-\varepsilon,2\sqrt{d-1}+\varepsilon]$ tends to $1$ when $n$ tends to infinity. In a colloquial language, this means that a random graph is close to be Ramanujan. Observe that this result does not say anything about Ramanujan graphs, i.e. when $\varepsilon=0$. It is indeed challenging to construct explicit examples, as discussed for instance in [9, Introduction]. Our results give the following
Corollary 1.8
With the notation of Definition 1.2, if $p$ is congruent to 1 modulo 12, and $H$ contains $\ell$, and $\det H=(\mathbb{Z}/N\mathbb{Z})^{\times}$, then the isogeny graph $G(p,\ell,H)$ is a Ramanujan graph.
The first three conditions guarantee that the adjacency matrix is symmetric, see Proposition 2.2.2; if we drop them, our main results say that the graphs are Ramanujan in some generalized sense. Corollary 1.8 can be applied for instance to isogeny graphs with Borel level structure.
With the same spirit, people have looked at expander sequences of graph. A sequence of $d$-regular connected finite graphs $G_{i}$ is an expander sequence if the adjacency matrices $A_{i}$ are symmetric, the number of vertexes tends to infinity, and there exists a constant $\varepsilon>0$ independent of $i$ such that the spectral gap of $G_{i}$ is at least $\varepsilon$ for every $i$. We again refer to [15, 29] and references therein for a detailed discussion. Observe that in loc. cit. the definition is given in terms of the expansion constant; our definition in terms of spectral gap is equivalent to the classical one because of the Cheeger inequality ([29, Sections 4.4 and 4.5] and [15, Section 1.2]). The importance of constructing explicit examples is highlighted for instance in [29, Section 2.1] (and our examples are explicit in the sense of Definition 2.3 of loc. cit.) or [32]. The following Corollary provides many new examples of expander sequences of graphs.
Corollary 1.9
Fix a prime $\ell$ and a sequence of graphs $\{G_{i}\}=\{G(p_{i},\ell,H_{i})\}$ with $p_{i}\equiv 1\pmod{12}$ and $H_{i}<\mathrm{GL}_{2}(\mathbb{Z}/N_{i}\mathbb{Z})$ a subgroup containing $\ell$, with determinant $\det H_{i}=(\mathbb{Z}/N_{i}\mathbb{Z})^{\times}$, and such that $[\mathrm{GL}_{2}(\mathbb{Z}/N_{i}\mathbb{Z}):H_{i}]\cdot p_{i}$ tends to infinity.
Then $\{G_{i}\}$ is an expander sequence of graphs.
The first example where Corollary 1.9 can be applied is the classical sequence of isogeny graphs: $N_{i}=1$ for every $i$, and $p_{i}$ grows. New examples are for instance when $p_{i}$ is fixed and $[\mathrm{GL}_{2}(\mathbb{Z}/N_{i}\mathbb{Z}):H_{i}]\to\infty$, which happens e.g. if $N_{i}$ grows, and $H_{i}$ is of a fixed type such as Borel or Cartan; or when $p_{i}$ grows, $N_{i}$ and $H_{i}$ can be anything.
Again, if we drop the condition of $p_{i}$ congruent to 1 modulo 12, and $H_{i}$ containing $\ell$, the adjacency matrix is not longer symmetric and the sequence is expander in a generalized sense.
Let us now look at the distribution of all eigenvalues, the bulk of the spectrum following the terminology of [29, Section 7.1], see also [45, Section 8]. Given a sequence $G_{i}$ of graphs as above and an angle $\theta$, we introduces the probability measure
$\mu(G_{i},\theta):=\frac{1}{|\sigma(A_{i},\theta)|}\sum_{\lambda\in\sigma(A_{i},\theta)}\delta_{\lambda}\,,$
where $\sigma(A_{i},\theta)$ is the set of eigenvalues of the adjacency matrix $A_{i}$ with phase $\theta$ or $\theta+\pi$,and $\delta_{\lambda}$ is a Dirac mass at $\lambda$; of course the definition makes sense only if $|\sigma(A_{i},\theta)|\neq 0$ If all eigenvalues of $A_{i}$ are real, we omit the dependence from $\theta$. Varying $\theta$, the limits of the sequences $\{\mu(G_{i},\theta)\}$, if they exist, gives the asymptotic distribution of the spectrum of $G_{i}$. Let us also introduce the Kesten–McKay measure (also known as Kesten–McKay law or distribution)
$\mu_{\ell}:=\frac{\ell+1}{\pi}\frac{\sqrt{\ell-x^{2}/4}}{\ell(\ell^{1/2}+\ell^{-1/2})^{2}-x^{2}}dx$ (1.10)
supported in the Hasse interval $[-2\sqrt{\ell},2\sqrt{\ell}]$; it is the asymptotic distribution of the eigenvalues of a random sequence of $\ell+1$-regular graphs with increasing number of vertexes, see [36], [29, Theorem 7.2] and references therein.
The following result, which relies on the theory of modular forms, is a corollary of Theorems 6.5.5 and 6.7.1.
Corollary 1.11
Fix a subgroup $H<\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$, a prime number $\ell$ coprime with $N$, and let $\{p_{i}\}$ be an increasing sequence of prime numbers not dividing $N\ell$. Let $G_{i}=G(p_{i},\ell,H)$,
  • If $H=\{\mathrm{Id}\}$, i.e. $G_{i}$ are isogeny graphs with full level structure, given $k^{\prime}$ the order of $\ell$ in $(\mathbb{Z}/N\mathbb{Z})^{\times}$, then for every $\theta$ in $\mathbb{Z}\frac{\pi}{k^{\prime}}$ we have
    $\lim_{i\to\infty}\mu(G_{i},\theta)=e^{i\theta}\mu_{\ell}$
    and for all other choices of $\theta$ there are no eigenvalues.
  • If $H$ is the Borel, then all eigenvalues are real and
    $\lim_{i\to\infty}\mu(G_{i})=\mu_{\ell}$
  • If $H=\{\left(\begin{smallmatrix}*&0\\ *&1\\ \end{smallmatrix}\right)\}$, i.e. the $G_{i}$’s are graphs with torsion point structure, denoting $k^{\prime}$ the order of $\ell$ in $(\mathbb{Z}/N\mathbb{Z})^{\times}$, then for every $\theta$ in $\mathbb{Z}\frac{\pi}{k^{\prime}}$ we have
    $\lim_{i\to\infty}\mu(G_{i},\theta)=e^{i\theta}\mu_{\ell}$
    and for all other choices of $\theta$ there are no eigenvalues.
  • If $H$ is a non-split Cartan, then all eigenvalues are real and
    $\lim_{i\to\infty}\mu(G_{i})=\mu_{\ell}$
It is instructive to note that Corollary 1.11 alone does not imply that all eigenvalues are contained in the Hasse interval: it does not prevent a small number of eigenvalues to lie outside the support of the asymptotic distribution.
By general graph theory, Corollary 1.11 implies that $G_{i}$ has few cycles, more precisely the number of cycles of a fixed length divided by the number of vertexes of $G_{i}$ tends to zero when $i$ tends to infinity, see [36] and [45, Theorem 10].
1.3 Relation with isogeny based cryptography
Usually the security, and sometime also the design, of protocols from isogeny based cryptography relies on features of isogeny graphs. Often the security is related to the mixing time, the number of cycles, or to the spectral gap of the graphs. All these features can be studied looking at the spectrum of the adjacency matrix. (We again refer to [29] or other textbooks in Graph Theory or Markov Chains for a general discussion of this topic).
The first appearance of isogeny graphs in cryptography is the Charles-Lauter hash function [12], where the digest of a message is computed trough a random walk on a classical isogeny graph.
Anoher important istance of isogeny based cryptography is the key exchange protocol SIDH [24]. In this protocol, the public key is two vertexes on the isogeny graph with full level structure at a known distance, and the private key is a walk between them. This protocol has been broken around a year ago [11, 35, 44]: if $N$ is big enough with respect to the length of the walk, as in SIDH, there are efficient algorithms to find a path between the two vertexes. If $N$ is small with respect to the length of the walk, still we do not known an efficient algorithm to find such a path. Observe that, by general graph theory, the difficulty of finding such a walk can be related to the spectral gap and mixing time of the graph. By now, many variant of SIDH have been proposed. Public keys can always been interpreted as a pair of vertexes on an isogeny graph with convenient level structure. Depending on the protocol, their distance can be either a public or a private parameter. Private keys are a walks between the two vertexes. For instance, in [22], the group $H$ defining the level structure is the group of scalar matrices; in [8], the authors look at the group of circulant matrices. It is not known if there is some intrinsic property of the isogney graphs which makes the path finding problem more difficult for some level structure rather than others.
From a different perspective, in [7] a Zero Knowledge Proof is defined using random walks on the isogeny graph with Borel level structure. A precise analysis of the spectral gap and, consequently, of the mixing time, is used to prove that the Proof of Knowledge is statistically secure.
Isogeny based cryptography is an active area of research, of course here we do not attempt to make a comprehensive review. We hope that our work can support its development.
1.4 Relation with other works
The Ramanujan property of classical isogeny graphs, i.e. without level structure, is usually attributed to A. K. Pizer [41]. In loc. cit. there is a sketch of the proof, which builds on previous work by Brandt, Eichler and Deligne. The approach is different from ours, as it goes directly trough modular forms. The main idea is to use the so called Brandt pairing to relate elliptic curves to modular forms, and eventually use the theory of Hecke operators and results similar to our Theorem 3.8. This approach is taken up in full details in [7, Section 3], where it is extended to the case of isogeny graph with Borel level structure. The relation with modular forms is also discussed in our Section 6.
An approach similar to ours is suggested in [43] and [21], however in these papers isogeny graphs are not the main focus. Building on [43], [33] studies the zeta-function of isogeny graphs with Borel level structure.
Isogeny graphs of ordinary curves are studied by Kohel [31], they have a rather different (and simpler!) structure than the supersingular ones, sometime they go by the names of volcano graphs or jellyfish graphs
The Borel level structure case is also studied by Arpin in [3]. Other interesting papers are [4, 2]. It is worth pointing out that in [28, 5], there is nice bound on the number of cycles on classical isogeny graphs obtained using different from ours. A few weeks after our this article appeared on ArXiv, Page and Wesolowski in [40] gave an independent proof of a variant of Theorem 1.6 using Jacquet–Langlands correspondence. In [34] graphs with full level structures are studied as infinite towers, studying the Galois and the connectedness properties.
Acknowledgments
We have had the pleasure and the benefit of conversations about the topics of this paper with S. Arpin, A. Basso, P. Caputo, L. De Feo, T. Morrison, M. Sala, M. Salvi, R. Schoof, S. Vigogna and F. Viviani. The first author also would like to thank the organizers and the partecipants of the Banff/Bristol 2023 workshop “Isogeny graphs in Cryptography” for many discussions on the topics of this paper.
Both authors are supported by the MIUR Excellence Department Project MatMod@TOV awarded to the Department of Mathematics, University of Rome Tor Vergata, the “National Group for Algebraic and Geometric Structures, and their Applications” (GNSAGA - INdAM), and the PRIN PNRR 2022 “Mathematical Primitives for Post Quantum Digital Signatures”. The second author is also supported by “Programma Operativo Nazionale (PON) “Ricerca e Innovazione” 2014-2020.
2 First properties of isogeny graphs
We fix $p,\ell,N,H$ as in Definition 1.2, together with the isogeny graph $G=G(p,\ell,H)$ and its vertices $V$. The adjacency matrix $A$ defines a linear operator $A\colon\mathbb{C}^{V}\to\mathbb{C}^{V}$ which maps a vertex $v$ to $\sum v_{i}$, where the sum runs over all edges $v\to v_{i}$ coming out of $v$.
2.1 Automorphisms of isogeny graphs
For the next subsection, and for other reasons later on, we will need the following operators.
Definition 2.1.1 (Diamond and matricial automorphisms)
Let $G$ be as in Definition 1.2. For every $g$ in the normalizer $N(H)$ of $H$ in $\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$ we define an automorphism
$\langle g\rangle\colon G\to G$
$(E,\phi)\mapsto(E,\phi\circ g)$
In particular, for every $d$ in $(\mathbb{Z}/N\mathbb{Z})^{\times}$, the diamond operator $\langle d\rangle$ is automorphism associated to the diagonal matrix $\left(\begin{smallmatrix}d&\,\\ \,&d\\ \end{smallmatrix}\right)$.
Observe that if $d=\left(\begin{smallmatrix}d&\,\\ \,&d\\ \end{smallmatrix}\right)$ belongs to $H$, then $\langle d\rangle$ is the identity. Moreover, even if $-1\notin H$, then $\langle-1\rangle$ is the identity because $(E,-\phi)$ is always isomorphic to $(E,\phi)$.
Proposition 2.1
For every $p$, $\ell$, $N$, and $H$, the isogeny graph $G(p,\ell,H)$ is the quotient of the isogeny graph with full level structure $G(p,\ell,\{\mathrm{Id}\})$ by the action of $H$ given in Definition 2.1.1. In particular, the spectrum of the adjacency matrix of $G(p,\ell,H)$ is a subset of the spectrum of the adjacency matrix of $G(p,\ell,\{\mathrm{Id}\})$.
Using Proposition 2.1, one could deduce most of our results from the case of full level structure. However we have preferred to give proofs that directly work for any level structure.
Another construction that yields automorphisms of isogeny graphs is the following.
Definition 2.1.2 (Frobenius automorphism)
Let $\sigma$ be the Frobenius of $\overline{\mathbb{F}_{p}}/\mathbb{F}_{p}$, then
$\langle\sigma\rangle\colon G\to G$
maps a vertex $(E,\phi)$ to the conjugated $(E^{\sigma},\phi^{\sigma}:=\sigma\circ\phi)$, and an isogeny to the conjugated by $\sigma$.
Notice that up to isomorphism, we can suppose that each elliptic curve $E_{i}$ in our graph is defined over $\mathbb{F}_{p^{2}}$ and that the Frobenius $\mathrm{Frob}_{p^{2}}\colon E_{i}\to E_{i}$ acts as $[-p]$. Since the map $\sigma\colon E(\overline{\mathbb{F}_{p}})\to E^{\sigma}(\overline{\mathbb{F}_{p}})$ coincides with the action of $\mathrm{Frob}_{p}\colon E\to E^{\sigma}$, we deduce that $\langle\sigma\rangle^{2}=\langle p\rangle$ on the graph: indeed, for each vertex $(E_{i},\phi_{i})$ we have
$\langle\sigma\rangle^{2}(E_{i},\phi_{i})=(E_{i}^{\sigma^{2}},\sigma^{2}\circ\phi_{i})=(E_{i},\mathrm{Frob}_{p^{2}}\circ\phi_{i})=(E_{i},[-p]\circ\phi_{i})=\langle-p\rangle(E_{i},\phi_{i})=\langle p\rangle(E_{i},\phi_{i})\,,$
where the last equality is true because $\langle-1\rangle$ is the identity.
Further automorphisms will be introduced in Definition 3.11.
2.2 Hermitian form and diagonalization
With keep the notation of Definition 1.2. We introduce the followin hermitian form $H$ on $\mathbb{C}^{V}$
$H\left((E_{i},\phi_{i}),(E_{j},\phi_{j})\right)=\delta_{ij}a_{i}\,,$ (2.2.1)2.2.1( 2.2.1 )
with $a_{j}=|\mathrm{Aut}(E_{i},\phi_{i})|$ and $\delta_{ij}$ is the Kronecker delta.
Proposition 2.2.2 (Adjoint of the adjacency matrix)
Let $G$ and $A$ be as in Definition 1.2 and let $A^{*}$ be its adjoint with respect to the Hermitian form ((2.2.1)2.2.1( 2.2.1 )). Then,
$A^{*}=\langle\ell^{-1}\rangle A\,.$
The adjacency matrix $A$ is diagonalizable, and the angles of its eigenvalues lie in $\mathbb{Z}\tfrac{\pi}{k^{\prime}}$, where $k^{\prime}$ is the minumum positive integer such that $\ell^{k^{\prime}}\mathrm{Id}\in H$. In particular:
  • the operators $A$ and $A^{*}$ commute, are both diagonalizable, have the same spectrum, and hence are conjugated.
  • if $\ell$ belongs to $H$, then $A=A^{*}$ and the spectrum of $A$ is real;
  • if $\ell$ belongs to $H$ and $p$ is congruent to 1 modulo 12 and $\ell$ belong to $H$, the adjacency matrix is symmetric.
Proof.
For the first part, we need to prove that, given vertices $(E_{i},\phi_{i})$ and $(E_{j},\phi_{j})$ we have
$H\big{(}A{\cdot}(E_{i},\phi),(E_{j},\phi_{j})\big{)}=H\big{(}(E_{i},\phi_{i}),\langle\ell^{-1}\rangle A{\cdot}(E_{j},\phi_{j})\big{)}\,,$ (2.2.3)2.2.3( 2.2.3 )
where we interpret $(E_{i},\phi_{i})$ and $(E_{j},\phi_{j})$ as elements of $\mathbb{C}^{V}$. Let $L$ be the set of degree $\ell$ morphisms $(E_{i},\phi_{i})\to(E_{j},\phi_{j})$, and let $M$ be the set of degree $\ell$ morphisms $(E_{j},\phi_{j})\to(E_{i},[\ell]\phi_{i})$. Then, using the definition of $A$, and the definition ((2.2.1)2.2.1( 2.2.1 )) of $H$, we find that
$H(A{\cdot}(E_{i},\phi),(E_{j},\phi_{j}))=\frac{\#L\#\mathrm{Aut}(E_{j},\phi_{j})}{\#\mathrm{Aut}(E_{j},\phi_{j})}\,,\quad H((E_{i},\phi_{i}),\langle\ell\rangle A{\cdot}(E_{j},\phi_{j}))=\frac{\#M\cdot\#\mathrm{Aut}(E_{i},\phi_{i})}{\#\mathrm{Aut}(E_{i},[\ell]\phi_{i})}\,.$
We notice that $\mathrm{Aut}(E_{i},\ell\phi_{i})$ equals $\mathrm{Aut}(E_{i},\phi_{i})$ as subgroup of $\mathrm{Aut}(E_{i})$. Hence equation ((2.2.3)2.2.3( 2.2.3 )) is equivalent to the fact that $L$ and $M$ have the same cardinality: indeed duality of isogenies gives a bijection between the two.
Since diamond operators commute with $A$, then $A$ is a normal operator, hence diagonalizable. Moreover, the adjoint of $A^{k^{\prime}}$ is equal to $\langle\ell^{k^{\prime}}\rangle A^{k^{\prime}}=A^{k^{\prime}}$, hence $A^{k^{\prime}}$ is Hermitian and has real eigenvalues. We deduce that for each $\lambda$ in the spectrum of $A$, its power $\lambda^{k^{\prime}}$ is real, hence the angle of $\lambda$ lies in $\mathbb{Z}\tfrac{\pi}{k^{\prime}}$.
The operator $A^{*}$ is also diagonalizable. Since $A$ and $A^{*}$ commute, they have the same eigenvectors. The corresponding eigenvalues are conjugated. Since $A$ is real, its specturm is invariant under conjugation, hence $A$ and $A^{*}$ have the same spectrum.
If $p$ is congruent to 1 modulo 12, all supersingular elliptic curves have {±$\{\pm 1\}$ as automorphism group, and hence all vertexes $(E_{i},\phi_{i})$ have the same number $a_{i}$ of automorphisms: if $-1\in H$, then $a_{i}=2$, otherwise $a_{i}=1$. Then the Hermitian form from Equation ((2.2.1)2.2.1( 2.2.1 )) is a multiple of the standard form, and being self-adjoint coincides with being symmetric. ∎
Remark 2.2.4
Since the Hermitian form ((2.2.1)2.2.1( 2.2.1 )) is presented in diagonal form, it is easy to write down the entries of $A^{*}$: for each $i$ we have
$A^{*}((E_{i},\phi_{i}))=a_{i}\sum_{j}a_{j}^{-1}(E_{j},\phi_{j})\,,$ (2.2.5)2.2.5( 2.2.5 )
where $a_{i},a_{j}$ are as in Equation ((2.2.1)2.2.1( 2.2.1 )), and the sum runs over all edges $(E_{j},\phi_{j})\to(E_{i},\phi_{i})$, namely all the edges in $G$ with end-point $(E_{i},\phi_{i})$. We notice that the entries of $A^{*}$ are integers: any vertex $(E_{j},\phi_{j})$ appearing in the right hand side of ((2.2.5)2.2.5( 2.2.5 )) has multiplicity $a_{i}a_{j}^{-1}\cdot(\#S/a_{i})=\#S/a_{j}$, for $S$ the set of degree $\ell$ isogenies $(E_{j},\phi_{j})\to(E_{i},\phi_{i})$; since $\mathrm{Aut}(E_{j},\phi_{j})$ acts freely on $S$ by precomposition, then #$\#S/a_{j}$ is an integer.
2.3 Weil pairing and reduction of the main theorems to Theorem 5.7
To formulate the following arguments we introduce the oriented Caley graph $C=C(N,\det H,\ell)$: vertexes are the element of $R_{H}=\mu_{N}^{\times}(\overline{\mathbb{F}}_{p})/\det H$, there is an edge from $\xi_{1}$ to $\xi_{2}$ if and only if $\xi_{2}=\xi_{1}^{\ell}$. (In the Borel level structure case, this graph is just one vertex with no edges.)
Since $\mu_{N}^{\times}(\overline{\mathbb{F}}_{p})$ is a principal homogeneous space for the right action of $(\mathbb{Z}/N\mathbb{Z})^{\times}$, the graphs $C(N,\det H,\ell)$ has simple structure: it is the disjoint union of $n$ cycles $C_{1},\ldots C_{n}$, each having the form of a loop:
{tikzpicture}
with $k$ the order of $\ell$ in $(\mathbb{Z}/N\mathbb{Z})^{\times}/\det H$ and $n=\phi(N)/(k|\det H|)$. In particular, the adjacency matrix $P_{i}$ of each $C_{i}$ is the cyclic permutation matrix on $k$ elements; its spectrum is thus the set $\mu_{k}(\mathbb{C})$ of the $k$-th roots of unity in $\mathbb{C}$.
If two elliptic curves with level structure are connected by a degree $\ell$ isogeny, then [46, Chapter III, Proposition 8.2] implies that Weil invariant of the level structures are one the $\ell$-th power of the other, hence we have the following result.
Proposition 2.3.1
The Weil invariant (see Definition 1.5) of a level structure gives a surjective map of graphs
$w\colon G(p,\ell,H)\to C(N,\det H,\ell)\,.$ (2.3.2)2.3.2( 2.3.2 )
Moreover, in the language of Definitions 2.1.1 and 2.1.2, we have $w(\langle g\rangle(E,\phi))=w((E,\phi))^{\det(g)}$ and $w(\sigma(E,\phi))=w((E,\phi))^{p}$
An example of isogeny graph together with the map $w$ is given in Figure 1.
Denoting $V(\cdot)$ the set of vertexes of a graph, the above map extends to a linear map
$w_{*}\colon\mathbb{C}^{V(G)}\longrightarrow\mathbb{C}^{V(C)}\,.$ (2.3.3)2.3.3( 2.3.3 )
Fix a connected components $C_{i}$ of $C$, let $G_{i}:=w^{-1}(C_{i})$ (this definition of $G_{i}$ coincides with the one in the Introduction), then the above map restricts to a morphism
$w_{i,*}\colon\mathbb{C}^{V(G_{i})}\longrightarrow\mathbb{C}^{V(C_{i})}\,.$ (2.3.4)2.3.4( 2.3.4 )
Remark 2.2 (Description of $\operatorname{Ker}(w_{i,*})$)
The kernel of the map $w_{i,*}$ defined in Equation ((2.3.4)2.3.4( 2.3.4 )) will play an important role in this paper, so let us describe it explicitly.
In the the Borel level structure case, $\operatorname{Ker}(w_{i,*})$ is the space of linear combination of vertexes of the graphs whose coefficient sum up to zero, i.e. the orthogonal complement of the vector $(1,\dots,1)$ in $\mathbb{C}^{V(G)}$ for the standard scalar product.
In general, $\operatorname{Ker}(w_{i,*})$ is equal to the subspace of $\mathbb{C}^{V(G_{i})}$ spanned by linear combinations of elements of $V(G_{i})$ with the same Weil invariant and such that the coefficients sum up to zero. In other words, calling $V_{\xi}$ the set of vertices of $G$ with Weil invariant $\xi\in R_{H}$, then
$\ker(w_{*})=\bigoplus_{\xi\in R_{H}}\left\{x\in\mathbb{C}^{V_{\xi}}:\sum x_{v}=0\right\}\,,\quad\ker(w_{i,*})=\bigoplus_{\xi\in V(C_{i})}\left\{x\in\mathbb{C}^{V_{\xi}}:\sum x_{v}=0\right\}$
where we identify $\mathbb{C}^{V(G)}$ with the direct sum of the various $\mathbb{C}^{V_{\xi}}$.
Since $w$ is a map of graphs, $\operatorname{Ker}(w_{i,*})$ is stable for the action of the adjacency matrix $A_{i}$ of $G_{i}$.
Proposition 2.3.5
Let $G$ be an isogeny graph as in Definition 1.2, $G_{i}$ be one of its subgraphs defined above, and $A_{i}$ the adjacency matrix of $G_{i}$.
The spectrum of $A_{i}$ is equal to the union of $(\ell+1)\mu_{k}(\mathbb{C})$ and the spectrum of $A_{i}$ restricted to $\operatorname{Ker}(w_{i,*})$, where $\mu_{k}(\mathbb{C})$ is the group of $k$-th roots of unity in $\mathbb{C}$.
Proof.
Let $\xi_{j}$, for $j=1,\dots,k$ be the vertexes of $C_{i}$. Let $v_{j}$ in $\mathbb{C}^{V(G_{i})}$ be the sum of all elliptic curves with Weil invariant $\xi_{j}$. Let $U$ be the span of $v_{1},\dots,v_{k}$. Then $U$ is a complement of $\operatorname{Ker}(w_{i,*})$ in $\mathbb{C}^{V(G_{i})}$, and, as $A_{i}$-module, it is isomorphic to the quotient $\mathbb{C}^{V(G_{i})}/\operatorname{Ker}(w_{i,*})$. Using the relation $(\ell{+}1)P_{i}\circ w_{i,*}=w_{i,*}\circ A_{i}$, we see that the restriction of $A_{i}$ to $U$ is conjugated to the matrix
$\left(\begin{matrix}0&&&\\ \vdots&&(\ell{+}1)\mathrm{Id}_{k-1}&\\ 0&&&\\ (\ell{+}1)&0&\cdots&0\end{matrix}\right)$
hence it is diagonalizable with spectrum $(\ell{+}1)\mu_{k+1}(\mathbb{C})$. ∎
The study of the spectrum of $A_{i}$ to $\operatorname{Ker}(w_{i,*})$ is rather delicate; Sections 3, 4 and 5 are devoted to the proof of the following result.
Theorem 2.3.6 (= Theorem 5.7)
Let $G$ be as in Definition 1.2, let $G_{i}$ be one of its subgraphs defined above, with adjacency matrix $A_{i}$, acting on the kernel $\operatorname{Ker}(w_{i,*})$ of the map ((2.3.4)2.3.4( 2.3.4 )). Then the modules of the eigenvalues of $A_{i}$ restricted to $\operatorname{Ker}(w_{i,*})$ are contained in the open Hasse interval $(-2\sqrt{\ell},2\sqrt{\ell})$.
Let us now take this result for granted and deduce the theorems stated in the Introduction.
Corollary 2.3.7
With the notation as in Theorem 2.3.6, each $G_{i}$ is connected. If $p,\ell$ and $\det N(H)$ generate $(\mathbb{Z}/N\mathbb{Z})^{\times}$, then all $G_{i}$’s are isomorphic.
Proof.
By general graph theory ([15, Proposition 1.1.2]), the number of connected component of an $\ell{+}1$ regular graph is the multiplicity of the eigenvalues $\ell{+}1$ for the adjacency matrix, hence Proposition 2.3.5 and Theorem 2.3.6 implies that $G_{i}$ is connected.
For the second part we notice that $p,\ell$ and $\det N(H)$ generate $(\mathbb{Z}/N\mathbb{Z})^{\times}$ if and only if $\langle p,\det N(H)\rangle$ acts transitively on the set of orbits $\{C_{1},\ldots,C_{n}\}$. If, for $g$ in $N(H)$, $\det(g)$ maps $C_{i}$ to $C_{j}$, then $\langle g\rangle$ and $\langle g^{-1}\rangle$ give an isomorphism between $G_{i}$ and $G_{j}$. Analogously, if $p$ maps $C_{i}$ to $C_{j}$, then $\langle\sigma\rangle$ gives an isomorphism between $G_{i}$ and $G_{j}$. ∎
[…]
[width=16cm]Grafo23bis.png
Figure 1: This is an example of isogeny graph $G(p,\ell,H)$ with $p=23$, $\ell=3$ and $H=\langle\left(\begin{smallmatrix}5&6\\ 2&1\\ \end{smallmatrix}\right),\left(\begin{smallmatrix}1&2\\ 0&1\\ \end{smallmatrix}\right),\left(\begin{smallmatrix}7&0\\ 2&7\\ \end{smallmatrix}\right),\left(\begin{smallmatrix}5&0\\ 0&5\\ \end{smallmatrix}\right),\left(\begin{smallmatrix}2&7\\ 7&1\\ \end{smallmatrix}\right),\left(\begin{smallmatrix}1&4\\ 0&1\\ \end{smallmatrix}\right),\left(\begin{smallmatrix}1&0\\ 4&1\\ \end{smallmatrix}\right)\rangle$ the only index 8 subgroup of $\mathrm{GL}_{2}(\mathbb{Z}/8\mathbb{Z})$.
       Color indicates the elliptic curve, or equivalently the $j$-invariant. The level structure is given through a matrix: for each of the three elliptic curves we have chosen a (non-canonical) basis of the $8$-torsion, and the matrix gives the change of basis. For mere visual clarity, arrows going up are black, arrows going down are gray.
       The graph $G(p,\ell,H)$ has two components, $G_{1}$ and $G_{2}$, which correspond via the Weil invariant to the two connected components $C_{1}$ and $C_{2}$ of the Cayley graph, depicted on the right. Since each $C_{i}$ has two vertexes, each $G_{i}$ is bipartite, see Remark 1.7. At the bottom, the graph without level structure.
Proof of Theorems 1.4 and 1.6 The statement about the connected components is Corollary 2.3.7. Diagonalizability and the angles of the eigenvalues are given in Proposition 2.2.2. The eigenvalues of absolute value $\ell{+}1$ are described in Proposition 2.3.5. The size of the other eigenvalues is described in Theorem 2.3.6.
2.4 Isomorphism between Borel and Cartan level structure
Here we define an isomorphism between graphs with Borel and Cartan level structure; because of this isomoprhism, in the sequel, when will spell out our results in special cases, we will consider only Borel level structures and not the Cartan ones.
Fix $p$ and $\ell$ distinct primes; let $N$ be a positive integer coprime with $p$ and $\ell$; let $B_{0}(N^{2})$ be the Borel subgroup of $\mathrm{GL}_{2}(\mathbb{Z}/N^{2}\mathbb{Z})$ and $T(N)$ the split Cartan of $\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$. Consider the maps
$F\colon G(p,\ell,B_{0}(N^{2}))$ $\to G(p,\ell,T(N))$ (2.3)
$(E,C)$ $\mapsto(E/NC,C/NC,E[N]/NC)$
Proposition 2.4
The map $F$ defined in Equation (2.3) gives an isomorphism of graphs.
3 Preliminary results on modular curves
Given a scheme $S$, a generalized elliptic curve $\pi\colon E\to S$ is a family of genus one curve whose singular members are Néron polygons, see [17, Chapter II].
Given a positive integer $N$ and a subgroup $H$ of $\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$ we denote $\mathcal{M}_{H}$ the stack over $\mathbb{Z}[1/N]$ parametrizing generalized elliptic curves with level $H$ structure, namely generalized elliptic curves $\pi\colon E\to S$ such that the fibers are either smooth or Néron polygons with $N$ edges, together with an isomorphism $\phi$ of the $N$ torsion of $E$ with $(\mathbb{Z}/N\mathbb{Z})^{\oplus 2}_{S}$; two level structures $\phi_{1}$ and $\phi_{2}$ are isomorphic if étale locally on $S$ there exists an $h$ in $H$ such that $\phi_{1}=\phi_{2}\circ h$. The stack $\mathcal{M}_{H}$ is a proper and smooth Deligne-Mumford stack over $\mathbb{Z}[1/N]$, see [17, Section IV.3, and Theorem 3.4].
For the proofs of our results, we need a more general definition of level structure. The most general notion is the one of Drinfeld level structure, see [30]; in this paper we will need only need a generalization of Borel level structure already discussed in [17], which we recall below.
For every positive integer $k$, let $B_{0}(k)=\{\left(\begin{smallmatrix}*&0\\ *&*\\ \end{smallmatrix}\right)\}$ be the standard Borel subgroup of $\mathrm{GL}_{2}(\mathbb{Z}/k\mathbb{Z})$. Given $M=Nq_{1}\cdots q_{r}$, with $q_{i}$ prime powers that are pairwise coprime and prime and to $N$, we consider level structure associated to subgroups $K$ of $\mathrm{GL}_{2}(\mathbb{Z}/M\mathbb{Z})$ of the form
$K=H\times B_{0}(q_{1})\times\cdots\times B_{0}(q_{r})\quad<\quad\mathrm{GL}_{2}(\mathbb{Z}/M\mathbb{Z})=\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})\times\prod_{i=1}^{r}\mathrm{GL}_{2}(\mathbb{Z}/q_{i}\mathbb{Z})\,,$ (3.1)
for $H$ a subgroup of $\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$. When $r=1$ and $q_{1}=p$ is prime, we write
$H_{p};=H\times B_{0}(p)\quad<\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})\times\mathrm{GL}_{2}(\mathbb{Z}/p\mathbb{Z})=\mathrm{GL}_{2}(\mathbb{Z}/Np\mathbb{Z})\,.$ (3.2)
For these kind of subgroups, a level $K$ structure on a generalized elliptic curve $\pi\colon E\to S$ is the datum of a level $H$ structure, and, for each $q_{i}$, a cyclic locally free group subscheme $G_{i}$ of rank $q_{i}$ which intersect every irreducible component of every geometric fiber of $\pi$.
Since a Borel subgroup $B_{0}(q)$ is the stabilizer of a line in $(\mathbb{Z}/q\mathbb{Z})^{2}$, we observe that over $\mathbb{Z}[1/M]$ this second definition of $K$ level structure is equivalent to the previous one, i.e. to an isomorphism between the $M$-torsion and $(\mathbb{Z}/M\mathbb{Z})^{\oplus 2}$ up to the action of $K$. If, for example, $q_{i}$ is prime, over $q_{i}$ the kernel of the Frobenius is a “new” example of $B_{0}(q_{i})$-structure for an elliptic curve/$\overline{F_{q_{i}}}$ which does not fit in the previous definition.
The stack $\mathcal{M}_{K}$ parametrizes generalized elliptic curves such that the Néron polygons have only $M$ edges with level $K$ structure; it is a proper and regular Deligne-Mumford stack over $\mathbb{Z}[1/N]$, it smooth outside the $\overline{\mathbb{F}}_{q_{i}}$ points parametrizing supersingular elliptic curves, see [17, Chapter V, Theorem 1.6, Propisition 1.10, Variants 1.14 and 1.20].
For $d$ in $(\mathbb{Z}/N\mathbb{Z})^{\times}$, the diamond operator $\langle d\rangle$ is an automorphism of $\mathcal{M}_{K}$: we let
$\langle d\rangle(E,\phi,G_{1},\dots,G_{r}):=(E,d\phi,G_{1},\dots,G_{r})\,.$ (3.3)
We now introduce two key maps, that will play more than one role for us
$\mathrm{pr}_{p}\colon\mathcal{M}_{H_{p}}\to\mathcal{M}_{H}\,,\quad$ $\mathrm{pr}_{p}(E\to S,\phi,C)=(E\to S,\phi)\,,$ (3.4)
$\mathrm{quot}_{p}\colon\mathcal{M}_{H_{p}}\to\mathcal{M}_{H}\quad$ $\mathrm{quot}_{p}(E\to S,\phi,C)=(E/C\to S,\pi_{C}\circ\phi)\,,$
where $\pi_{C}$ is the quotient map $E\to E_{C}$
Following [17, Section V], we first use them to study the fiber $\mathcal{M}_{H_{p},\mathbb{F}_{p}}=\mathcal{M}_{H_{p}}\times\operatorname{Spec}\mathbb{F}_{p}$. The maps $\mathrm{pr}_{p}$ and $\mathrm{quot}_{p}$ have right inverse when restricted to $\mathcal{M}_{H_{p},\mathbb{F}_{p}}$. Indeed, an elliptic curve $E$ over $\overline{\mathbb{F}}_{p}$ has only two subgroup or rank $p$: the kernel of the Frobenius and the kernel of the Verschiebung (which, by definition, is the dual isogeny of the Frobenius). They are equal if and only if the curve is supersingular. We obtain two morphisms
$\mathrm{pr}_{p,p}^{-1}\colon\mathcal{M}_{H,\mathbb{F}_{p}}$ $\longrightarrow\mathcal{M}_{H_{p},\mathbb{F}_{p}}\,,\qquad(E/S/\mathbb{F}_{p},\phi)\mapsto(E/S/\mathbb{F}_{p},\phi,\ker(\mathrm{Frob}))\,,$ (3.5)
$\mathrm{quot}_{p,p}^{-1}\colon\mathcal{M}_{H,\mathbb{F}_{p}}$ $\longrightarrow\mathcal{M}_{H_{p},\mathbb{F}_{p}}\,,\qquad(E/S/\mathbb{F}_{p},\phi)\mapsto(E^{(p)}/S/\mathbb{F}_{p},\phi\circ(\cdot\tfrac{1}{p})\circ\mathrm{Frob},\ker(\mathrm{Ver}))\,,$
which provide a description of $\mathcal{M}_{H_{p},\mathbb{F}_{p}}$ as the union of two copies of $\mathcal{M}_{H,\mathbb{F}_{p}}$ nodally attached at the supersingular elliptic curves, see [17, Section 5, Theorem 1.16 and Variant 1.18]. Here we apologize for an abuse of notations: $\mathrm{pr}_{p,p}^{-1}$ and $\mathrm{quot}_{p,p}^{-1}$ are not the inverse of $\mathrm{pr}_{p,p}=\mathrm{pr}_{p,\mathbb{F}_{p}}$ and $\mathrm{quot}_{p,p}=\mathrm{quot}_{p,\mathbb{F}_{p}}$, but just the right inverse.
Every Deligne-Mumford stack $\mathcal{M}$ admits a coarse space $M$, in particular $\mathcal{M}_{K}$ has a coarse space $M_{K}$. Every map between stacks, such as $\mathrm{pr}_{p}$ and $\mathrm{quot}_{p}$, induces a map between coarse spaces. A key fact is that in our set-up the formation of the coarse space is compatible with base change. More precisely, let $\ell$ be any prime number not dividing $N$ (possibly it can also be a divisor of the $q_{i}$’s); the universal property of coarse spaces gives a map from the coarse space of $\mathcal{M}_{K,\mathbb{F}_{\ell}}$ to $M_{K,\mathbb{F}_{\ell}}:=M_{K}\times\mathbb{F}_{\ell}$. In [17, Cor 6.10 page 145] it is shown that this map is an isomorphism (observe that if $\ell$ divides $N$ then this compatibility is not known for general $H$, see for instance [30, Section 8.5]).
We also use the maps (3.4) to define the Hecke operator $T_{\ell}$.
Definition 3.6 (Hecke operators)
With $K$ as in Equation (3.1), and for a prime $\ell$ which does not divide $M$, the Hecke operator $T_{\ell}$ is the map
$T_{\ell}:=(\mathrm{quot}_{\ell})_{*}\circ\mathrm{pr}_{\ell}^{*}\colon\mathrm{Pic}(\mathcal{M}_{K}/\mathbb{Z}[1/N])\to\mathrm{Pic}(\mathcal{M}_{K}/\mathbb{Z}[1/N])\,,$
where the push-forward is a cycle push-forward.
The analogue definition works for the coarse space $M_{K}$.
Observe that the diamond operator $\langle d\rangle$, which is defined for every $d$ which does not divide $N$, commutes with $\mathrm{pr}_{\ell}$, $\mathrm{quot}_{\ell}$ and $T_{\ell}$. From the description of the curves $\mathcal{M}_{K_{\ell},\mathbb{F}_{\ell}}$ we also obtain the following celebrated description of the restriction of the Hecke operator $T_{\ell}$ to $\mathrm{Pic}^{0}\mathcal{M}_{K,\mathbb{F}_{\ell}}$.
Theorem 3.7 (Eichler-Shimura relation)
With the notations of Definition 3.6, denoting by $T_{\ell,\mathbb{F}_{\ell}}$ the restriction of $T_{\ell}$ to either $\mathrm{Pic}^{0}(\mathcal{M}_{K,\mathbb{F}_{\ell}})$ or $\mathrm{Pic}^{0}(M_{K,\mathbb{F}_{\ell}})$, we have
$T_{\ell,\mathbb{F}_{\ell}}=\mathrm{Frob}_{*}+\langle\ell\rangle_{*}\mathrm{Frob}^{*}$
where $\langle\ell\rangle$ is the diamond automorphism (3.3) and $\mathrm{Frob}$ is the Frobenius of the curve $\mathcal{M}_{K,\mathbb{F}_{\ell}}$ or $M_{K,\mathbb{F}_{\ell}}$.
Proof.
We first prove the result of the stacks. Looking at the description of $\mathrm{quot}_{\ell,\mathbb{F}_{\ell}}$ and $\mathrm{pr}_{\ell,\mathbb{F}_{\ell}}$ on the two irreducible components of $\mathcal{M}_{K_{\ell}}$, we can write
$T_{\ell,\mathbb{F}_{\ell}}=(\mathrm{quot}_{\ell,\mathbb{F}_{\ell}}\circ\mathrm{pr}_{\ell,\mathbb{F}_{\ell}}^{-1})_{*}\circ(\mathrm{pr}_{\ell,\mathbb{F}_{\ell}}\circ\mathrm{pr}_{\ell,\mathbb{F}_{\ell}}^{-1})^{*}+(\mathrm{quot}_{\ell,\mathbb{F}_{\ell}}\circ\mathrm{quot}_{\ell,\mathbb{F}_{\ell}}^{-1})_{*}\circ(\mathrm{pr}_{\ell,\mathbb{F}_{\ell}}\circ\mathrm{quot}_{\ell,\mathbb{F}_{\ell}}^{-1})^{*}$
Both $\mathrm{pr}_{\ell,\mathbb{F}_{\ell}}\circ\mathrm{pr}_{\ell,\mathbb{F}_{\ell}}^{-1}$ and $\mathrm{quot}_{\ell,\mathbb{F}_{\ell}}\circ\mathrm{quot}_{\ell,\mathbb{F}_{\ell}}^{-1}$ are the identity on $\mathrm{Pic}^{0}\mathcal{M}_{K,\mathbb{F}_{\ell}}$, so we are left with
$T_{\ell,\mathbb{F}_{\ell}}=(\mathrm{quot}_{\ell,\mathbb{F}_{\ell}}\circ\mathrm{pr}_{\ell,\ell}^{-1})_{*}+(\mathrm{pr}_{\ell,\mathbb{F}_{\ell}}\circ\mathrm{quot}_{\ell,\ell}^{-1})^{*}$
We observe that $(\mathrm{quot}_{\ell,\mathbb{F}_{\ell}}\circ\mathrm{pr}_{\ell,\ell}^{-1})_{*}=\mathrm{Frob}_{*}$ because it maps $(E,\phi)$ to $(E^{(\ell)},\mathrm{Frob}\circ\phi)$ To conclude, $(\mathrm{pr}_{\ell,\mathbb{F}_{\ell}}\circ\mathrm{quot}_{\ell,\ell}^{-1})^{*}=\langle\ell\rangle_{*}\mathrm{Frob}^{*}$ because it maps $(E,\phi)$ to $(E^{(\ell)},\mathrm{Frob}\circ\phi\circ\left(\cdot\frac{1}{\ell}\right))$.
The property on the coarse spaces follows from their universal property.
The spectral bounds in Theorem 2.3.6 will eventually be a consequence of the following bound, which in turn is a consequence of the above mentioned Eichler-Schimura relation and Weil’s conjecture.
Theorem 3.8 (Bound on the eigenvalues of the Hecke operator)
[…]
Proof.
[…]
The Frobenius $\mathrm{Frob}_{X}$ is the endomorphism $\mathrm{Frob}_{*}$ appearing in Theorem 3.7. The maps $\mathrm{Frob}_{*}$ and $\mathrm{Frob}^{*}$ commutes, $\mathrm{Frob}^{*}\circ\mathrm{Frob}_{*}$ is the multiplication by $\ell$, hence also $\mathrm{Frob}_{*}$ has eigenvalues of complex absolute value $\ell^{i/2}$. The map $\langle\ell\rangle$ is an automorphism of finite order of $X$, hence its eigenvalues are root of unity.
Since the maps $\mathrm{Frob}_{*}$, $\mathrm{Frob}^{*}$ and $\langle\ell\rangle$ commute, the claim follows from Theorem 3.7. ∎
We close this section by introducing some automorphisms of modular curves, mirroring and expanding the list in Section 2.1.
Definition 3.9 (Matricial automorphisms)
Given a level structure $K=H\times\prod B(q_{i})$ as in (3.1), for any element $g$ in the normalizer $N(H)<\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$ of $H$, the automorphism $\langle g\rangle\colon\mathcal{M}_{K}\to\mathcal{M}_{K}$ maps a curve $(E,\phi,G_{1},\ldots G_{r})$ to $(E,\phi\circ g,G_{1},\ldots G_{r})$.
In particular, for every $d$ in $(\mathbb{Z}/N\mathbb{Z})^{\times}$, the diamond operator $\langle d\rangle$ in (3.3) is the automorphism associated to the diagonal matrix $\left(\begin{smallmatrix}d&\,\\ \,&d\\ \end{smallmatrix}\right)$.
Definition 3.10 (Fricke automorphism)
For a level structure $H_{p}$, the Fricke automorphism $\sigma\colon\mathcal{M}_{H_{p}}\to\mathcal{M}_{H_{p}}$ maps a curve $(E,\phi,G)$ to $(E/G,\pi\circ\phi,E[p]/G)$, where $\pi\colon E\to E/G$ is the projection.
Slightly more in general, we give the following
Definition 3.11 (Atkin-Lehner automorphisms)
Given $K=H\times\prod B(q_{i})$ as in (3.1), each $q_{i}$ yields the Atkin-Lehner map
$w_{q_{i}}\colon\mathcal{M}_{K}\to\mathcal{M}_{K}\,,\quad(E,\phi,G_{1},\ldots G_{r})\mapsto\left(E/G_{i},\pi_{i}\circ\phi,\pi_{i}(G_{1}),\ldots,E[q_{i}]/G_{i},\ldots,\pi_{i}(G_{r})\right)$ (3.12)
where $\pi_{i}\colon E\to E/G_{i}$ is the projection. Given an isogeny graph of the form $G=G(p,\ell,K)$, its vertices are tuples $(E,\phi,G_{1},\ldots,G_{r})$, and formula (3.12) defines an automorphism of $G$.
4 Relation between modular curves and isogeny graphs
In this section we explain the relation between the isogeny graph, together with its the adjacency matrix, and the coarse moduli space $M_{H_{p},\mathbb{F}_{p}}$, together with the Hecke operator $T_{\ell}$. See Remark 4.10 for the analysis on the stack.
We fix $p,N,H$ as in Definition 1.2. The maps (3.4) give the desingularization
$\mathrm{pr}_{p,p}^{-1}\sqcup\mathrm{quot}_{p,p}^{-1}\colon M_{H,\mathbb{F}_{p}}\sqcup M_{H,\mathbb{F}_{p}}\to M_{H_{p},\mathbb{F}_{p}}\,.$ (4.1)
Since the singularities of $M_{H_{p},\mathbb{F}_{p}}$ are nodal, the pull-back induces an exact sequence
$0\to T\to\mathrm{Pic}^{0}\left(M_{H_{p},\mathbb{F}_{p}}\right)\longrightarrow\mathrm{Pic}^{0}\left(M_{H,\mathbb{F}_{p}}\right)^{\times 2}\to 0$ (4.2)
with $T$ the toric part of the semi-abelian variety $\mathrm{Pic}^{0}\left(M_{H_{p},\mathbb{F}_{p}}\right)$.
To understand $T$, we need first to count the connected components of $M_{H,\overline{\mathbb{F}}_{p}}$. To this end, recall that the Weil invariant of a level structure, see Definition 1.5, gives a morphism
$w\colon M_{H}\to\operatorname{Spec}\left(\mathbb{Z}[\tfrac{1}{N},\zeta_{N}]^{\det(H)}\right)$
where $\zeta_{N}$ is a primitive $N$-th root of the unity, see [17, Chapter 3, Subsection 3.20], and the exponentiation to $\det(H)$ means that we take invariants of $\det(H)\subset(\mathbb{Z}/N\mathbb{Z})^{\times}=\mathrm{Gal}(\mathbb{Q}(\zeta_{N})/\mathbb{Q})$. If we base change to a field of characteristic prime to $N$, the fibers of $w$ are irreducible, see [17, Chapter 3, Corollary 5.6]. In particular, there is a bijection between the connected components of $\mathcal{M}_{H,\overline{\mathbb{F}}_{p}}$ and $R_{H}=\mu_{N}^{\times}(\overline{\mathbb{F}}_{p})/\det H$, see Definition 1.5 and above. Call these components $M_{\xi}$, for $\xi$ in $R_{H}$. The discussion below Equation (3.5) implies that the map $\mathrm{pr}_{p,\mathbb{F}_{p}}$ is surjective and gives a bijection between the connected components of $M_{H,\overline{\mathbb{F}}_{p}}$ and the ones of $M_{H_{p},\overline{\mathbb{F}}_{p}}$.
By definition, points on $T$ correspond to line bundles $L$ over $M_{H_{p}}$ such that both $\left(\mathrm{pr}_{p,p}^{-1}\right)^{*}L$ and $\left(\mathrm{quot}_{p,p}^{-1}\right)^{*}L$ are trivial. As recalled in Appendix A, to describe such an $L$ we need to give a scalar for each node of $M_{H_{p},\mathbb{F}_{p}}$, modulo a diagonal action of $\mathbb{G}_{\mathrm{m}}$ for every connected component $M_{\xi}$. Recall that the nodes of $M_{H_{p},\mathbb{F}_{p}}$ are the points representing supersingular curves. Call $V_{\xi}$ the set of vertexes of $G=G(p,\ell,H)$ with Weil invariant $\xi$, which are in turn the points of $M_{\xi}$ such that $\mathrm{pr}_{p,p}^{-1}(v)$ is singular in $M_{H_{p},\overline{\mathbb{F}}_{p}}$. With this notation we have a canonical isomorphism
$T\cong\,\,\prod_{\xi\in R_{H}}T_{\xi}\quad\textrm{with}\quad T_{\xi}:=\mathbb{G}_{\mathrm{m}}^{V_{\xi}}/\mathbb{G}_{\mathrm{m}}\,.$ (4.3)
For the groups of characters $T^{\vee}:=\mathrm{Hom}(T,\mathbb{G}_{\mathrm{m}})$ and $T_{\xi}^{\vee}:=\mathrm{Hom}(T_{\xi},\mathbb{G}_{\mathrm{m}})$, we obtain
$T^{\vee}=\bigoplus_{\xi\in R_{H}}T_{\xi}^{\vee}\quad\textrm{with}\quad T_{\xi}^{\vee}\cong\left\{x\in\mathbb{Z}^{V_{\xi}}\,:\,\textstyle\sum_{v\in V_{\xi}}x_{v}=0\right\}\,.$ (4.4)
This identifies $T^{\vee}$ with a submodule of $\mathbb{Z}^{V}$, and comparing with Remark 2.2 we see that $T^{\xi}\otimes\mathbb{C}=\ker(w_{*})$ inside $\mathbb{C}^{V}$. Moreover, consider $R_{H}=C_{1}\sqcup\ldots\sqcup C_{n}$ the decomposition of $R_{H}$ into the orbits of $\xi\to\xi^{\ell}$, as in the discussion below Proposition 2.3.1: for each $C_{i}$, again comparing with Remark 2.2 we have
$\bigoplus_{\xi\in C_{i}}T^{\vee}_{\xi}\otimes\mathbb{C}=\left\{x\in\mathbb{C}^{V_{\xi}}\,:\,\textstyle\sum_{v\in V_{\xi}}x_{v}=0\right\}=\operatorname{Ker}(w_{i,*})\,,$ (4.5)
where $\operatorname{Ker}(w_{i,*})$ the subspace of $\mathbb{C}^{V}$ described in Remark 2.2.
Theorem 4.6
Let $G=G(p,\ell,H)$ be the graph in Definition 1.2, with $G_{i}$ the subgraphs defined above Theorem 1.6, and let $T=\prod_{\xi\in R_{H}}T_{\xi}$ be the maximal torus of $\mathrm{Pic}^{0}(\mathcal{M}_{H_{p},\overline{\mathbb{F}}_{p}})$, as in Equations (4.2) and (4.3).
For each $i$, the isomorphism (4.5) intertwines the action of the Hecke operator $T_{\ell}$ with the adjoint action of the adjency matrix of the graph $G_{i}$: i.e. the following diagram is commutative
[…]
where $\operatorname{Ker}(w_{i,*})$ is the subspace of $\mathbb{C}^{V}$ described in Remark 2.2, and $A^{*}$ is the adjoint of the adjacency matrix $A$ with respect to the Hermitian form ((2.2.3)2.2.3( 2.2.3 )), see also Proposition 2.2.2.
Proof.
Let $V$ be the set of vertices of $G$. Equation (4.4) gives an embedding of $T^{\vee}$ and $\bigoplus_{\xi\in C_{i}}T^{\vee}_{\xi}$ inside $\mathbb{Z}^{V}$, and Proposition A.7 tells us that $T_{\ell}\colon T^{\vee}\to T^{\vee}$ (and in particular also its restriction to $\bigoplus_{\xi\in C_{i}}T^{\vee}_{\xi}$) extends to a map $T_{\ell}\colon\mathbb{Z}^{V}\to\mathbb{Z}^{V}$. It is enough to prove the commutativity of the diagram
[…]
In particular it is enough checking the commutativity on the elements $(E_{i},\phi_{i})$ of the canonical basis of $\mathbb{Z}^{V}\otimes\mathbb{C}$. Since geometrically we have $T_{\ell}=(\mathrm{quot}_{\ell})_{*}\circ\mathrm{pr}_{\ell}^{*}$, then Proposition A.7 gives
$T_{\ell}(E_{i},\phi_{i})=\!\!\sum_{(E_{j},\phi_{j},C)}\!\!\mathrm{ord}_{(E_{j},\phi_{j},C)}(\mathrm{quot}_{\ell}){\cdot}\mathrm{pr}_{\ell}(E_{j},\phi_{j},C)=\!\!\sum_{(E_{j},\phi_{j},C)}\!\!\mathrm{ord}_{(E_{j},\phi_{j},C)}(\mathrm{quot}_{\ell})\cdot(E_{j},\phi_{j})$ (4.7)
where $(E_{j},\phi_{j},C)$ varies in the fiber $\mathrm{quot}_{\ell}^{-1}(E_{i},\phi_{i})\subset M_{H_{\ell}}(\overline{\mathbb{F}}_{p})$.
To compute the $\mathrm{ord}(\mathrm{quot}_{\ell})$ we start by noticing that when $H$ structures are rigid (i.e. when $\mathrm{Aut}(E,\phi)=\{1\}$ for each $(E,\phi)$ in $M_{H}(\overline{\mathbb{F}}_{p})$), then $\mathrm{ord}(\mathrm{quot}_{\ell})=1$: indeed $\mathrm{quot}_{\ell}$ has degree $\ell{+}1$ and duality of isogenies gives a bijection between the set of points $(E_{i},\phi_{j},C)\in\mathrm{quot}_{\ell}^{-1}(E_{i},\phi_{i})$ and the set of points $(E_{i},\tfrac{1}{\ell}\phi_{i},C)\in\mathcal{M}_{H_{\ell}}(\overline{\mathbb{F}}_{p})$ which has cardinality $\ell{+}1$ because $\mathrm{Aut}(E_{i},\phi_{i})$ is trivial, hence for different subgroups $C_{1},C_{2}\subset E_{i}[\ell]$ the triples $(E_{i},\tfrac{1}{\ell}\phi_{i},C_{2})$ and $(E_{i},\tfrac{1}{\ell}\phi_{i},C_{2})$ are not isomorphic.
For general $H$ structure, even not rigid, write $M_{H,\mathbb{F}_{p}}=M_{K,\mathbb{F}_{p}}/G$ for $K$ a rigid level structure and $G$ a finite group, with quotient map $\pi_{G}$ (for example take $K$ to be full-level structures of level $3N$, see [30, Corollary 4.7.2], and $G<\mathrm{GL}_{2}(\mathbb{Z}/3N\mathbb{Z})$ to be the inverse image of $H$ under reduction modulo $N$). Analogously we have $M_{H_{\ell}}=M_{K_{\ell}}/G$, with quotient map $\pi_{G,\ell}$. Now, given $(E_{j},\phi_{j},C)$ supersingular point on $\mathcal{M}_{H_{p,\ell}}$, we can lift it to a point $(E_{j},\psi_{j},C)$ on $M_{K_{p,\ell}}$, and, using the commutation $\mathrm{quot}_{\ell}\circ\pi_{G,\ell}=\pi_{G}\circ\mathrm{quot}_{\ell}$, we compute
$\mathrm{ord}_{(E_{j},\phi_{j},C)}\mathrm{quot}_{\ell}$ $=\frac{\mathrm{ord}_{(E_{j},\psi_{j},C)}(\mathrm{quot}_{\ell}\circ\pi_{G,\ell})}{\mathrm{ord}_{(E_{j},\psi_{j},C)}\pi_{G,\ell}}=\frac{\mathrm{ord}_{(E_{j},\psi_{j},C)}(\pi_{G}\circ\mathrm{quot}_{\ell})}{\mathrm{ord}_{(E_{j},\psi_{j},C)}\pi_{G,\ell}}$
$=\frac{\mathrm{ord}_{(E_{j},\psi_{j},C)}(\mathrm{quot}_{\ell})\cdot\mathrm{ord}_{(E_{i},\psi_{i})}\pi_{G}}{\mathrm{ord}_{(E_{j},\psi_{j},C)}\pi_{G,\ell}}=\frac{1\cdot|\mathrm{Aut}(E_{i},\phi_{i})|}{|\mathrm{Aut}(E_{j},\phi_{j},C)|}\,.$
Substituting in Equation (4.7), and using the definition of $a_{i}$ in ((2.2.1)2.2.1( 2.2.1 )), we get
$T_{\ell}(E_{i},\phi_{i})=\sum_{(E_{j},\phi_{j},C)}\frac{|\mathrm{Aut}(E_{i},\phi_{i})|}{|\mathrm{Aut}(E_{j},\phi_{j},C)|}\cdot(E_{j},\phi_{j})=a_{i}\sum_{(E_{j},\phi_{j},C)}|\mathrm{Aut}(E_{j},\phi_{j},C)|^{-1}\cdot(E_{j},\phi_{j})\,,$ (4.8)
where the sums run over the isomorphism classes of triples $(E_{j},\phi_{j},C)\in M_{H_{\ell}}(\overline{\mathbb{F}}_{p})$ such that $\mathrm{quot}_{\ell}(E_{j},\phi_{j},C):=(E_{j}/C,\pi_{C}\circ\phi_{j})$ is isomorphic to $(E_{i},\phi_{i})$. We want to compare the last term of Equation (4.8) with the description of $A^{*}$ given in Remark 2.2.4.
Observe that $(E_{j},\phi_{j})$ appears in the right hand side of (4.8) if and only if there is an arrow $(E_{j},\phi_{j})\to(E_{i},\phi_{i})$. The number of such arrows equals the number of nontrivial subgroups $C\subset E_{i}[\ell]$ such that $(E_{j}/C,\pi_{C}\circ\phi_{j})\cong(E_{i},\phi_{i})$. Two triples $(E_{j},\phi_{j},C_{1})$ and $(E_{j},\phi_{j},C_{2})$ give the same element of $M_{H_{\ell}}(\overline{\mathbb{F}}_{p})$ if and only if there exist $\sigma$ in $\mathrm{Aut}(E_{j},\phi_{j})/\mathrm{Aut}(E_{j},\phi_{j},C_{1})$. Such $\sigma$, if it exists, is unique because we have quotiented out exactly by the stabilizer of $(E_{j},\phi_{j},C_{1})$ in $\mathrm{Aut}(E_{j},\phi_{j})$. We conclude that the coefficient of $(E_{j},\phi_{j})$ in the right hand side of Equation (4.8) is
$a_{i}\sum_{\begin{subarray}{c}0\subsetneq C\subsetneq E_{j}[\ell]\text{ s.t.}\\ (E_{j}/C,\pi_{C}\circ\phi_{j})\cong(E_{i},\phi_{i})\end{subarray}}|\mathrm{Aut}(E_{j},\phi_{j})/\mathrm{Aut}(E_{j},\phi_{j},C)|^{-1}|\mathrm{Aut}(E_{j},\phi_{j},C)|^{-1}$
As, in Remark 2.2.4 we have $a_{i}=|\mathrm{Aut}(E_{j},\phi_{j})|$, we have the claim.
The following proposition discuss the equivariance of the canonical isomorphism (4.5) with respect to the automorphisms of the graph and the modular curve.
Proposition 4.9
Keep the notation of Theorem 4.6. The canonical isomorphism of $T_{\ell}$-modules $A^{*}$-modules
[…]
also intertwines the Galois map (Definition 2.1.2) with the Fricke map 3.10 and it is equivariant with respect to matricial automorphisms coming from the normalizer of $H$ (Definitions 2.1.1 and 3.9) and, if there, to the Atkin-Lehner maps of $H$-structures (Definition 3.11), where automorphisms of modular curves act on the Picard groups, hence on $T^{\vee}$, via pull-back.
Proof.
This is an application of Proposition A.7 in the case where $G$ is the identity of $M_{H_{p},\mathbb{F}_{p}}$ and $F$ is one of the automorphisms of $M_{H_{p},\mathbb{F}_{p}}$ we have considered. In particular it is enough checking that the action of matricial automorphisms, respectively Atkin-Lehner automorphisms and Fricke map, on the supersingular points of $M_{H_{p},\mathbb{F}_{p}}$ is exactly the action of the corresponding automorphisms of the graph: in the first two cases this is straight forward; for the Fricke map we observe that, given a point $(E,\phi,\ker(\mathrm{Frob}_{p}))$ of $M_{H_{p},\mathbb{F}_{p}}(\overline{\mathbb{F}}_{p})$ representing a supersingular elliptic curve, we have
$\sigma\big{(}E,\phi,\ker(\mathrm{Frob}_{p}))=\big{(}E/\ker(\mathrm{Frob}_{p}),\pi\circ\phi,E[p]/\ker(\mathrm{Frob}_{p})\big{)}\,,$
which is equal to $\big{(}E^{\sigma},\sigma\circ\phi,\ker\mathrm{Frob}_{p}\big{)}$ because $E/\ker(\mathrm{Frob}_{p})$ is supersingular, hence $E[p]/\ker(\mathrm{Frob}_{p})$ must be equal to the kernel of its Frobenius, and moreover the quotient map $\pi\colon E\to E/\ker(\mathrm{Frob}_{p})$ is exactly the Frobenius map $\mathrm{Frob}_{p}\colon E\to E^{\sigma}$. We conclude that the Fricke map acts as the Galois map on $\overline{\mathbb{F}_{p}}$-points of $M_{H_{p},\mathbb{F}_{p}}$ representing supersingular elliptic curves. ∎
Remark 4.10 (Analogous construction on the moduli stack)
One could carry out the constructions of this section on the stack $\mathcal{M}_{H_{p},\mathbb{F}_{p}}$ rather than the coarse space $M_{H_{p},\mathbb{F}_{p}}$. Observe that when $p\geq 5$, so the characteristic of the base field does not divide the automorphism group, this stack is a twisted curve, as in [1, Section 2]. Twisted curve are also called stacky curves in the literature. At least in these cases, in loc. cit. is explained how the Picard group is an extensions of the Picard group of the coarse space by a finite étale group over $\mathbb{F}_{p}$ related to the automorphism groups. The study of this extension might give further information about isogeny graphs.
5 Proof of Theorem 2.3.6
Definition 5.1
Given $p,N,H$ as in Definition 1.2, let $\mathcal{A}=\mathcal{A}_{H,p}$ over $\mathbb{Z}[1/N]$ be the connected component of the identity of the kernel of the map
$\left(\mathrm{pr}_{p,*},\mathrm{quot}_{p,*}\right)\colon\mathrm{Pic}^{0}\left(M_{H_{p}}\right)\longrightarrow\mathrm{Pic}^{0}\left(M_{H}\right)\times\mathrm{Pic}^{0}\left(M_{H}\right)$
The action of the Hecke operator $T_{\ell}$, and the automorphism from Definitions 3.10, 3.9 and 3.11 preserve $\mathcal{A}$, hence we can and do consider their restriction to $\mathcal{A}$.
Proposition 5.2
Fix $p,N,H$ as in Definition 1.2. The fiber $\mathcal{A}_{\mathbb{F}_{p}}$ is equal to the torus $T$ introduced in Equation (4.2).
Proof.
Since $\mathrm{Pic}^{0}\left(M_{H,\mathbb{F}_{p}}\right)$ is an abelian variety, and there are no non-trivial map from a torus to an abelian variety, we have the inclusion $T\subseteq\mathcal{A}_{\mathbb{F}_{p}}$.
Since $\dim T=\dim\mathrm{Pic}^{0}\left(M_{H_{p,\overline{F}_{p}}}\right)-\dim\left(\mathrm{Pic}^{0}\left(M_{H_{\overline{F}_{p}}}\right)\times\mathrm{Pic}^{0}\left(M_{H_{\overline{F}_{p}}}\right)\right)$, to conclude we have to show that the reduction modulo $p$ of $\left(\mathrm{pr}_{p,*},\mathrm{quot}_{p,*}\right)$ is surjective.
We look at the resolution given by Equation (4.1) and we consider the map
$\lambda\colon\mathrm{Pic}^{0}(M_{H,\mathbb{F}_{p}})^{\times 2}\longrightarrow\mathrm{Pic}^{0}(M_{H_{p},\mathbb{F}_{p}})\,,\quad(x,y)\longmapsto(\mathrm{pr}_{p,p}^{-1})_{*}(x)+(\mathrm{quot}_{p,p}^{-1})_{*}(y)\,.$
By the same arguments used in the proof of Theorem 3.7, (or see also the diagram in [17, page 145]), we have that $\left(\mathrm{pr}_{p,*},\mathrm{quot}_{p,*}\right)_{\mathbb{F}_{p}}\circ\lambda$ equals $\left(\begin{smallmatrix}\mathrm{Id}&\mathrm{Frob}\\ \mathrm{Frob}&\mathrm{Id}\\ \end{smallmatrix}\right)$ as endomorphism of $\mathrm{Pic}^{0}(M_{H,\mathbb{F}_{p}})^{\times 2}$; this endomorphism is surjective, hence the same is true for $\left(\mathrm{pr}_{p,*},\mathrm{quot}_{p,*}\right)_{\mathbb{F}_{p}}$.
The following key technical lemma uses the theory of Néron models.
Lemma 5.3
Fix $p,N,H$ as in Definition 1.2 and let $\mathcal{A}=\mathcal{A}_{H_{p}}$. Then, for every endomorphism $F$ of $\mathcal{A}$ and every prime number $q$ not dividing $N$, we have
$\dim(\mathrm{Im}\left(F|_{\mathcal{A}_{\mathbb{C}}}\right))=\dim\left(\mathrm{Im}\left(F|_{\mathcal{A}_{\overline{\mathbb{F}_{q}}}}\right)\right)\,.$
Proof.
By [17, Proposition 6.7 and Theorem 6.9, pages 143-145], both $M_{H}/\mathbb{Z}[\tfrac{1}{N}]$ and $M_{H_{p}}/\mathbb{Z}[\tfrac{1}{N}]$ have reduced fibers, and geometrically irreducible generic fiber. Again by loc. cit., $M_{H}$ is regular, but $M_{H_{p}}$ might not be: it is smooth away from supersingular elliptic curves $(E,\phi,C)$ in characteristic $p$, and locally around such points it is isomorphic to $\mathbb{Z}_{p}[[w,z]]/(wz-p^{k})$, where $k$ is either #$\#Aut(E,\phi,C)$, or half of it if $-1$ is an automorphism. To reduce ourselves to the regular case we can blow-up the non-regular points. In this way, we introduce a chain of $\mathbb{P}^{1}$’s on the fiber over $p$; this chain does not alter the $\mathrm{Pic}^{0}$, hence we can assume by abuse of notation that also $M_{H_{p}}$ is regular.
We now localize at $q$ and apply [10, Theorem 4 (b), Section 9.5, page 267]: both $\mathrm{Pic}^{0}\left(M_{H_{p}}\right)$ and $\mathrm{Pic}^{0}\left(M_{H}\right)$ are the connected component of the identity of the Néron models of $\mathrm{Pic}^{0}\left(M_{H_{p}}\right)_{\mathbb{Q}}$ and $\mathrm{Pic}^{0}\left(M_{H}\right)_{\mathbb{Q}}$, hence $\mathcal{A}$ is the connected component of the identity of the Néron model of $\mathcal{A}_{\mathbb{Q}}$ (this last assertion can checked using the universal property of Néron models). Moreover, by Lemma 5.4 and [17, Proposition 6.7, page 143], $\mathcal{A}$ has semi-abelian reduction.
When there is semi-abelian reduction, by [10, Proposition 3, section 7.5, page 186], taking Néron models is exact up to isogeny, so we have the claim.
Lemma 5.4
Fix $p,\ell,H$ as in Definition 1.2. There is a (non-canonical) isomorphism of $T_{\ell}$ modules
[…]
where $T^{\vee}$ is the group of characters of the torus $T$ introduced in Equation (4.2) and $\mathcal{A}$ is the abelian variety in Definition 5.1. This isomorphism is also equivariant for the automorphism from Definitions 3.9, 3.10 and 3.11
Proof.
First we show that there exists a non-canonical isomorphism $\gamma$ of $T_{\ell}$-modules. For this it is enough showing a $\mathbb{Q}$-linear isomorphism between $T^{\vee}\otimes\mathbb{Q}$ and $H^{1,\text{sing}}(\mathcal{A}_{\mathbb{C}},\mathbb{Z})\otimes\mathbb{Q}$, as $\mathbb{Q}[x]$-modules, with $x$ acting as $T_{\ell}$. Since $\mathbb{Q}[x]$ is a PID, it is enough showing that for every polynomial $q$ in $\mathbb{Z}[x]$, the rank of $F:=q(T_{\ell})$ is equal on both spaces. The morphism $F$ is an endomorphism of $\mathcal{A}$. The rank of $F$ restricted to $T^{\vee}\otimes\mathbb{Q}$ is equal to $\dim(\mathrm{Im}\left(F|_{\mathcal{A}_{\overline{\mathbb{F}_{\ell}}}}\right))$. The rank of $F$ on $H^{1,\text{sing}}(\mathcal{A}_{\mathbb{C}},\mathbb{Q})=H^{1,\text{sing}}(\mathcal{A}_{\mathbb{C}},\mathbb{Z})\otimes\mathbb{Q}$ is equal to twice $\dim(\mathrm{Im}\left(F|_{\mathcal{A}_{\mathbb{C}}}\right))$. We obtain the claim by Lemma 5.3.
Now we have to show that we can choose a $\gamma$ which is equivariant for all automorphisms. Let $G$ be the group formed by these automorphisms. Theorem 4.6 and Proposition 2.2.2 imply that $T_{\ell}$ is semi-simple. Since $G$ commutes with $T_{\ell}$, then it preserves the eigenspaces of $T_{\ell}$. Each eigenspace is a $G$ module, and we have to show that these $G$ module are isomorphic. To this end, since $G$ is finite, it is enough to show that the characters are the same. This can be proved by looking at the rank of endomorphisms induced by polynomials in elements of $G$, and applying again Lemma 5.3.
The following lemma is a rather general fact
Lemma 5.5
Fix $p,\ell,H$ as in Definition 1.2. For any prime $\ell^{\prime}$ which does not divide $p\ell N$, we have an isomorphism of $T_{\ell}$ modules
[…]
Proof.
The isomorphism is given by the cospecialization map, let us explain the argument. By proper-smooth base change theorem (see [38, Theorem 20.4]), the cospecialization map
[…]
is an isomorphism. Since the cospecialization map is functorial, then it is an isomorphism of $T_{\ell}$ modules.
Moreover, since $\mathcal{A}_{\mathbb{C}}$ is a smooth variety over $\mathbb{C}$, then the comparison theorem [38, Theorem 21.1] tells us that, for each positive integer $k$, we have isomorphisms
[…]
Since the above isomorphism is functorial, then, again, it also an isomorphism of $T_{\ell}$ modules. The proof of the second statement is analogous. ∎
We are now ready to prove the main technical theorem about isogeny graphs used in Section 2.
Theorem 5.7 (= Theorem 2.3.6)
The modules of the eigenvalues of $A_{i}$ restricted to $\operatorname{Ker}(w_{i,*})$ are contained in the Hasse interval $(-2\sqrt{\ell},2\sqrt{\ell})$.
Proof.
To prove that the eigenvalues are contained in the closed Hasse interval $[-2\sqrt{\ell},2\sqrt{\ell}]$ Apply in the following order Theorem 4.6, Lemma 5.4, Lemma 5.5, and then the combination of Eichler-Shimura relation and Weil conjectures stated in Theorems 3.7, 3.8.
To show that the eigenvalues can not have absolute value equal to $2\sqrt{\ell}$ we need to pass trough the theory of modular forms. First we identify them with the eigenvalues of the Hecke operator on a space of modular forms using Theorem 6.5.5. We can now apply [13, Theorem 2.1].
6 Relation with modular forms
In this section we identify our spaces $\operatorname{Ker}(w_{i,*})$ from Remark 2.2 with spaces of modular forms. We start from the following lemma.
Lemma 6.1
Fix $p,\ell,H$ as in Definition 1.2. We have a (non-canonical) isomorphism of $T_{\ell}$ modules
$T^{\vee}\otimes\mathbb{C}\cong H^{0}(\mathcal{A}_{\mathbb{C}},\Omega^{1})$
where $T^{\vee}$ is the group of characters of the torus $T$ introduced in Equation (4.2) and $\mathcal{A}$ is the abelian variety in Definition 5.1. This isomorphism is equivariant for the automorphisms $u$ from Definitions 3.9, 3.10 and 3.11, acting by pullback on $\mathcal{A}$, hence as $u^{*,\vee}$ on $T^{\vee}$ and as $(u^{*})^{*}$ (see Remark 6.2) on the differentials of $\mathcal{A}_{\mathbb{C}}$.
Proof.
It is enough giving an isomorphism $T^{\vee}\otimes\mathbb{C}\cong H^{0}(\mathcal{A}_{\mathbb{C}},\Omega^{1})$, which is analogous to Lemma 5.4. ∎
Remark 6.2
For a map of curves $u\colon X\to Y$, we have the pullback $u^{*}\colon\mathrm{Pic}^{0}(Y)\to\mathrm{Pic}^{0}(X)$ and its pullback
$(u^{*})^{*}\colon H^{0}(\mathrm{Pic}^{0}(X),\Omega^{1})=H^{0}(X,\Omega^{1})\longrightarrow H^{0}(\mathrm{Pic}^{0}(Y),\Omega^{1})=H^{0}(Y,\Omega^{1})\,.$
Then, the above map is equal to the pushforward of differentials $u_{*}\colon H^{0}(X,\Omega^{1})\to H^{0}(Y,\Omega^{1})$. In particular, in Lemma 6.1, an automorphism $u$ acts as the restriction of $u_{*}$ on $H^{0}(\mathcal{A}_{\mathbb{C}},\Omega^{1})$
The above Lemma, together with Theorem 4.6, encourages the study differentials on $\mathcal{A}$: in Theorems 6.5.2 and 6.5.5 we relate these differentials with modular forms.
To include non-connected modular curves in our analysis, in the subsections 6.1 - 6.4 we recall the notation, mainly following [18], and collect some slightly cumbersome computations.
6.1 Complex points on modular curves
Analogously to [17, IV.5.3], using the definition $\mathbb{H}^{\pm}:=\mathbb{C}-\mathbb{R}$ and its “compactification” ¯±:=±$\overline{\mathbb{H}}^{\pm}:=\mathbb{H}^{\pm}\cup\mathbb{P}^{1}(\mathbb{Q})$, we have a (canonical) isomorphism of Riemann surfaces
$\begin{gathered}\text{GL}_{2}(\mathbb{Z})\backslash\big{(}\overline{\mathbb{H}}^{\pm}\times(\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})/H)\big{)}\overset{\sim}{\longrightarrow}M_{H}(\mathbb{C}),\end{gathered}$ (6.1.1)6.1.1( 6.1.1 )
where for each $\tau$’s in $\mathbb{H}^{\pm}$ (on proper elliptic curves) the map identifies
$(\tau,\gamma H)\longmapsto(E_{\tau},\phi_{\tau}\circ\gamma)=(\mathbb{C}/(\mathbb{Z}{+}\mathbb{Z}\tau),\phi_{\tau}\circ\gamma),\quad\phi_{\tau}\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)=\tfrac{\tau}{N}\,,\phi_{\tau}\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)=\tfrac{1}{N}\,.$
In the above isomorphism $\text{GL}_{2}(\mathbb{Z})$ acts by
$\begin{gathered}g\cdot(\tau,\gamma H):=(g(\tau),\bar{g}^{-T}\gamma H)\qquad\qquad\text{i.e.}\\ \left(\begin{smallmatrix}a&b\\ c&d\\ \end{smallmatrix}\right)(\tau,\gamma H)=\left(\tfrac{a\tau+b}{c\tau+d},\,\,\tfrac{1}{\det g}\left(\begin{smallmatrix}d&-c\\ -b&a\\ \end{smallmatrix}\right)\gamma H\right)\,.\end{gathered}$ (6.1.2)6.1.2( 6.1.2 )
For the subgroup $H_{p}<\mathrm{GL}_{2}(\mathbb{Z}/Np\mathbb{Z})$, Equation ((6.1.1)6.1.1( 6.1.1 )) can be rephrased as
$\Gamma^{0}(p)\backslash\big{(}\overline{\mathbb{H}}^{\pm}\times\tfrac{\mathrm{GL}_{2}(\mathbb{Z}/n\mathbb{Z})}{H}\big{)}$ $\overset{\sim}{\longrightarrow}M_{H_{p}}(\mathbb{C})\,,\qquad(\tau,\gamma)\longmapsto(E_{\tau},\phi_{\tau}\circ\gamma,\langle\tfrac{\tau}{p}\rangle)\,,$ (6.1.3)6.1.3( 6.1.3 )
where $\Gamma^{0}(p)$ is the subgroup of $\text{GL}_{2}(\mathbb{Z})$ made of matrices congruent to (*$\left(\begin{smallmatrix}*&0\\ *&*\\ \end{smallmatrix}\right)$ modulo $p$.
Using the above isomorphisms the maps $\mathrm{pr}_{p}$ and $\mathrm{quot}_{p}$ in (3.4) become
$\begin{gathered}\mathrm{pr}_{p},\,\mathrm{quot}_{p}\,\colon\Gamma^{0}(p)\backslash\big{(}\mathbb{H}^{\pm}\times\tfrac{\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})}{H}\big{)}\longrightarrow\text{GL}_{2}(\mathbb{Z})\backslash\big{(}\mathbb{H}^{\pm}\times\tfrac{\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})}{H}\big{)}\,,\\ \mathrm{pr}_{p}(\tau,\gamma)=(\tau,\gamma)\,,\qquad\mathrm{quot}_{p}(\tau,\gamma)=(\left(\begin{smallmatrix}1&0\\ 0&p\\ \end{smallmatrix}\right)\tau,\left(\begin{smallmatrix}p&0\\ 0&1\\ \end{smallmatrix}\right)\gamma)\end{gathered}$ (6.1.4)6.1.4( 6.1.4 )
The isomorphisms ((6.1.1)6.1.1( 6.1.1 )) ((6.1.3)6.1.3( 6.1.3 )) also help us recognize the components, over $\mathbb{C}$, of modular curves: choosing representatives $g_{1},\ldots,g_{r}$ for the quotient $\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})/(H\cdot\text{SL}_{2}(\mathbb{Z}/N\mathbb{Z}))$, we get the following (non-canonical) decomposition into connected components
$\begin{gathered}M_{H}(\mathbb{C})\cong\bigsqcup_{j=1}^{r}\Gamma_{g_{j}Hg_{i}^{-1}}\backslash\overline{\mathbb{H}}\,,\qquad\left(E_{\tau},\phi_{\tau}\circ g_{j}\right)\longmapsfrom(\tau,g_{j})\,,\\[-4.30554pt] M_{H_{p}}(\mathbb{C})\cong\bigsqcup_{j=1}^{r}\big{(}\Gamma^{0}(p){\cap}\Gamma_{g_{j}Hg_{i}^{-1}}\big{)}\backslash\overline{\mathbb{H}}\,,\qquad\left(E_{\tau},\phi_{\tau}\circ g_{j},\langle\tfrac{\tau}{p}\rangle\right)\longmapsfrom(\tau,g_{j})\,,\end{gathered}$ (6.1.5)6.1.5( 6.1.5 )
where ¯$\overline{\mathbb{H}}=\mathbb{H}\cup\mathbb{P}^{1}(\mathbb{Q})$ is the “compactification” of =$\mathbb{H}=\{\tau\in\mathbb{C}:\mathrm{Im}(\tau)>0\}$, and where
$\Gamma_{H}:=\{\gamma\in\text{SL}_{2}(\mathbb{Z}):\gamma^{T}\!\!\!\pmod{n}\text{ lies in }H\}.$
Remark 6.1.6
In Equation ((6.1.3)6.1.3( 6.1.3 )) we use $\Gamma^{0}(p)=\Gamma_{B^{0}(p)}$, with $B^{0}(p)$ the Borel group {(**0*)}$\ 0&*\\ \end{smallmatrix}\right)\}$ (notice the transposition in ((6.1.2)6.1.2( 6.1.2 ))). Since conjugation of the $H_{p}$ gives an isomorphic modular curve, we can also use $B_{0}(p)=\{\left(\begin{smallmatrix}*&0\\ *&*\\ \end{smallmatrix}\right)\}=\left(\begin{smallmatrix}0&1\\ 1&0\\ \end{smallmatrix}\right)B^{0}(p)\left(\begin{smallmatrix}0&1\\ 1&0\\ \end{smallmatrix}\right)^{-1}$, yielding a variant of ((6.1.3)6.1.3( 6.1.3 )):
$\Gamma_{0}(p)\backslash\big{(}\overline{\mathbb{H}}^{\pm}\times\tfrac{\mathrm{GL}_{2}(\mathbb{Z}/n\mathbb{Z})}{H}\big{)}$ $\overset{\sim}{\longrightarrow}M_{H_{p}}(\mathbb{C})\,,\qquad(E_{\tau},\phi_{\tau}\circ\gamma,\langle\tfrac{1}{p}\rangle)\longmapsfrom(\tau,\gamma)\,,$ (6.1.7)6.1.7( 6.1.7 )
for $\Gamma_{0}(p)=\Gamma_{B_{0}(p)}=\{\left(\begin{smallmatrix}a&b\\ c&d\\ \end{smallmatrix}\right)\in\mathrm{GL}_{2}(\mathbb{Z}):c\equiv 0\mod p\}$.
6.2 Modular forms and differentials
For any congruence subgroup $\Gamma$ of $\mathrm{SL}_{2}(\mathbb{Z})$, the map $f\mapsto f\mathrm{d}\tau$ gives an isomorphism between the space $S_{2}(\Gamma)$ of cuspidal modular forms of weight 2 and the space $H^{0}(\Gamma\backslash\overline{\mathbb{H}},\Omega^{1})$ of holomorfic differentials on $\Gamma\backslash\overline{\mathbb{H}}$, see [18, Section 3.3 and excercise 3.3.6] or [39, Theorem 2.3.2]. This, together with ((6.1.5)6.1.5( 6.1.5 )) implies the isomorphisms
$H^{0}(M_{H,\mathbb{C}},\Omega^{1})\cong\bigoplus_{j=1}^{r}S_{2}\left(\Gamma_{g_{j}Hg_{i}^{-1}}\right)\,,\qquad H^{0}(M_{H_{p},\mathbb{C}},\Omega^{1})\cong\bigoplus_{j=1}^{r}S_{2}\left(\Gamma_{g_{j}Hg_{i}^{-1}}\cap\Gamma^{0}(p)\right)$ (6.2.1)6.2.1( 6.2.1 )
6.3 Full level case
When $H=\{\mathrm{Id}\}$, we write $M_{N}$ for $M_{H}$ and $\Gamma(N)$ for $\Gamma_{H}$, which contains matrices in $\mathrm{SL}_{2}(\mathbb{Z})$ congruent to $\left(\begin{smallmatrix}1&0\\ 0&1\\ \end{smallmatrix}\right)$ modulo $N$. Choosing $\{g_{i}\}=\{\left(\begin{smallmatrix}a&0\\ 0&1\\ \end{smallmatrix}\right):a\in(\mathbb{Z}/N\mathbb{Z})^{\times}\}$, Equation ((6.1.5)6.1.5( 6.1.5 )) gives
$M_{N}(\mathbb{C})\cong\bigsqcup_{a\in(\mathbb{Z}/N\mathbb{Z})^{\times}}\!\!\!\!\Gamma(N)\backslash\overline{\mathbb{H}}\,,\quad M_{\{\mathrm{Id}\}\times B_{0}(p)}(\mathbb{C})\cong\bigsqcup_{a\in(\mathbb{Z}/N\mathbb{Z})^{\times}}\!\!\!\!\big{(}\Gamma^{0}(p){\cap}\Gamma(N)\big{)}\backslash\overline{\mathbb{H}}$ (6.3.1)6.3.1( 6.3.1 )
and, compatibly with this isomorphisms the map $\mathrm{pr}$, $\mathrm{quot}$ are
$\begin{gathered}\mathrm{pr}_{p},\,\mathrm{quot}_{p}\,\colon\!\!\!\!\bigsqcup_{a\in(\mathbb{Z}/N\mathbb{Z})^{\times}}\!\!\!\!\big{(}\Gamma^{0}(p){\cap}\Gamma(N)\big{)}\backslash\overline{\mathbb{H}}\longrightarrow\!\!\!\!\bigsqcup_{a\in(\mathbb{Z}/N\mathbb{Z})^{\times}}\!\!\!\!\Gamma(N)\backslash\overline{\mathbb{H}}\,,\\ \mathrm{pr}_{p}(\tau,a)=(\tau,a)\,,\qquad\mathrm{quot}_{p}(\tau,a)=(\left(\begin{smallmatrix}1&0\\ 0&q\\ \end{smallmatrix}\right)\tau,pa)\end{gathered}$ (6.3.2)6.3.2( 6.3.2 )
Moreover Equation ((6.2.1)6.2.1( 6.2.1 )) becomes
$H^{0}(M_{N,\mathbb{C}},\Omega^{1})\cong\bigoplus_{a\in(\mathbb{Z}/N\mathbb{Z})^{\times}}\!\!\!\!S_{2}\left(\Gamma(N)\right)=S_{2}\left(\Gamma(N)\right)\otimes_{\mathbb{C}}\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}\,,$ (6.3.3)6.3.3( 6.3.3 )
$H^{0}(M_{\{\mathrm{Id}\}\times B^{0}(p),\mathbb{C}},\Omega^{1})\cong\bigoplus_{a\in(\mathbb{Z}/N\mathbb{Z})^{\times}}\!\!\!\!S_{2}\left(\Gamma(N){\cap}\Gamma^{0}(p)\right)=S_{2}\left(\Gamma(N){\cap}\Gamma^{0}(p)\right)\otimes_{\mathbb{C}}\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}$
6.4 Hecke operators
As in [18, Section 5.1], we recall the definition of double coset operators: given $\Gamma_{1},\Gamma_{2}<\mathrm{SL}_{2}(\mathbb{Z})$ congruence subgroups, and given $\alpha\in\mathrm{GL}_{2}^{\det>0}(\mathbb{Q})$, we have the operator
$[\Gamma_{1}\alpha\Gamma_{2}]_{2}\colon M_{2}(\Gamma_{1})\to M_{2}(\Gamma_{2})\,,\quad f[\Gamma_{1}\alpha\Gamma_{2}]_{2}=\sum_{j}f[\alpha\gamma_{j}]_{2}\,,$ (6.4.1)6.4.1( 6.4.1 )
where $f[\left(\begin{smallmatrix}a&b\\ c&d\\ \end{smallmatrix}\right)]_{2}(\tau)=\det\left(\begin{smallmatrix}a&b\\ c&d\\ \end{smallmatrix}\right)\tfrac{1}{(c\tau+d)^{2}}f\left(\tfrac{a\tau+b}{c\tau+d}\right)$, and $\{\gamma_{j}\}$ is a set of representatives for $\Gamma_{3}\backslash\Gamma_{2}$, with $\Gamma_{3}=\alpha^{-1}\Gamma_{1}\alpha\cap\Gamma_{2}$. We can interpret the operator ((6.4.1)6.4.1( 6.4.1 )) as follows: we have maps
[…]
and, under the isomorphism ((6.2.1)6.2.1( 6.2.1 )), we have $[\Gamma_{1}\alpha\Gamma_{2}]_{2}=\pi_{2,*}\circ(\pi_{1}\alpha)^{*}$. A particular case are the classical Hecke operators in the theory of modular forms, see [18, Section 5.2]:
$\¯\tilde{T}_{\ell}:=[\Gamma\left(\begin{smallmatrix}1&0\\ 0&\ell\\ \end{smallmatrix}\right)\Gamma]_{2}=\pi_{*}\circ{\left(\begin{smallmatrix}1&0\\ 0&\ell\\ \end{smallmatrix}\right)}^{*}\,,\qquad{\left(\begin{smallmatrix}1&0\\ 0&\ell\\ \end{smallmatrix}\right)}\,,\pi\colon\left(\Gamma^{0}(\ell){\cap}\Gamma\right)\backslash\overline{\mathbb{H}}\to\Gamma\backslash\overline{\mathbb{H}}$ (6.4.3)6.4.3( 6.4.3 )
where ${\left(\begin{smallmatrix}1&0\\ 0&\ell\\ \end{smallmatrix}\right)}\tau=\tfrac{\tau}{\ell}$, $\pi\tau=\tau$, and we consider $\Gamma=\Gamma_{H}$ for $H<\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$ any subgroup that is normalized by diagonal matrices.
In the case $\Gamma=\Gamma(N)$, we want to compare $\tilde{T}_{\ell}$ with the Hecke operator $T_{\ell}$ in Definition 3.6. Indeed $T_{\ell}$ acts as $\mathrm{quot}_{\ell,*}\circ\mathrm{pr}_{\ell}^{*}$ on $\mathrm{Pic}^{0}(M_{N})$, hence it acts by pull back as $\mathrm{pr}_{\ell,*}\circ\mathrm{quot}_{\ell}^{*}$ on $H^{0}(\mathrm{Pic}^{0}(M_{N,\mathbb{C}}),\Omega^{1})=H^{0}(M_{N,\mathbb{C}},\Omega^{1})$. By ((6.3.3)6.3.3( 6.3.3 )), this space of differentials is isomorphic to $S_{2}\left(\Gamma(N)\right)\otimes_{\mathbb{C}}\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}$ and, under this identification, Equation ((6.3.2)6.3.2( 6.3.2 )) tells that that $\mathrm{pr}_{\ell,*}=\pi_{*}\otimes\mathrm{Id}$ and that $\mathrm{quot}_{\ell}^{*}={\left(\begin{smallmatrix}1&0\\ 0&\ell\\ \end{smallmatrix}\right)}^{*}\otimes\sigma_{\ell}$, where $\sigma_{\ell}\colon\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}\to\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}$ is the “shift by $\ell$” namely $(z_{a})\mapsto(z_{a\ell})$, and the maps $\pi,\left(\begin{smallmatrix}1&0\\ 0&\ell\\ \end{smallmatrix}\right)$ are the same appearing in ((6.4.3)6.4.3( 6.4.3 )). We deduce that
$T_{\ell}=\tilde{T}_{\ell}\otimes_{\mathbb{C}}\,\sigma_{\ell}\quad\text{ in }H^{0}(\mathrm{Pic}^{0}(M_{N,\mathbb{C}}),\Omega^{1})=S_{2}(\Gamma(N))\otimes_{\mathbb{C}}\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}\,.$ (6.4.4)6.4.4( 6.4.4 )
We have an analogous equality for $H=\{\mathrm{Id}\}\times B_{0}(p)$: using the second line in ((6.3.3)6.3.3( 6.3.3 ))
$T_{\ell}=\tilde{T}_{\ell}\otimes_{\mathbb{C}}\,\sigma_{\ell}\quad\text{ in }H^{0}(\mathrm{Pic}^{0}(M_{\{\mathrm{Id}\}\times B_{0}(p),\mathbb{C}}),\Omega^{1})=S_{2}(\Gamma^{0}(p){\cap}\Gamma(N))\otimes_{\mathbb{C}}\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}\,.$ (6.4.5)6.4.5( 6.4.5 )
6.5 Graphs versus modular forms
And now $H^{0}(\mathcal{A}_{\mathbb{C}},\Omega^{1})$: Definition 5.1 gives the canonical isomorphism
$H^{0}(\mathcal{A}_{\mathbb{C}},\Omega^{1})=\frac{H^{0}(M_{H_{p},\mathbb{C}},\Omega^{1})}{\mathrm{pr}_{p}^{*}\,H^{0}(M_{H,\mathbb{C}},\Omega^{1})+\mathrm{quot}_{p}^{*}\,H^{0}(M_{H_{p},\mathbb{C}},\Omega^{1})}\,.$
We start by looking at the case $H=\{\mathrm{Id}\}$, where Equation ((6.3.2)6.3.2( 6.3.2 )) gives an explicit description of $\mathrm{pr}_{p}^{*},\mathrm{quot}_{p}^{*}$. Instead of taking a quotient, we can take the orthogonal complement with respect to the Petersson inner product (see [18, Section 5.5]): following [43], we define the space of $p$-new forms as
$S_{2}^{p\text{-new}}(\Gamma^{0}(p){\cap}\Gamma(N)):=\Big{(}S_{2}(\Gamma(N))+S_{2}(\Gamma(N)[\left(\begin{smallmatrix}1&0\\ 0&p\\ \end{smallmatrix}\right)]_{2})\Big{)}^{\perp}\quad\subset S_{2}(\Gamma^{0}(p){\cap}\Gamma(N))\,,$
which, by the same arguments in [18, Proposition 5.5.2 and Proposition 5.6.2], is $\tilde{T}_{\ell}$-stable. In particular, using the description (((6.4.5)6.4.5( 6.4.5 )) of the Hecke operator, we get the isomorphism
$T_{\ell}\lefttorightarrow H^{0}(\mathcal{A}_{\{\mathrm{Id}\},p,\mathbb{C}},\Omega^{1})\cong S_{2}^{p\text{-new}}(\Gamma^{0}(p){\cap}\Gamma(N))\otimes_{\mathbb{C}}\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}\righttoleftarrow\tilde{T}_{\ell}\otimes\sigma_{\ell}$
To treat the case of a general $H$, we recall that $\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$ acts on $M_{\{\mathrm{Id}\}\times B_{0}(p)}$ by the law $(E,\phi,C)^{g}=(E,\phi\circ g,C)$. Using ((6.1.2)6.1.2( 6.1.2 )) and ((6.3.1)6.3.1( 6.3.1 )), we can characterise this action as follows:
[…]
where $\tilde{g_{a}}$ is any matrix in $\Gamma^{0}(p)$ that is congruent to $\left(\left(\begin{smallmatrix}a&0\\ 0&1\\ \end{smallmatrix}\right)g\left(\begin{smallmatrix}a&0\\ 0&1\\ \end{smallmatrix}\right)^{-1}\right)^{t}$ modulo $N$. We get an action of $\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$ by pullback on $H^{0}(\mathcal{A}_{\{\mathrm{Id}\},p,\mathbb{C}},\Omega^{1})\subset H^{0}(M_{\{\mathrm{Id}\}\times B_{0}(p)})$ as follows:
$\begin{gathered}\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})\lefttorightarrow S_{2}^{p\text{-new}}(\Gamma^{0}(p){\cap}\Gamma(N))\otimes_{\mathbb{C}}\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}=\bigoplus_{a\in(\mathbb{Z}/N\mathbb{Z})^{\times}}\!\!\!\!S_{2}^{p\text{-new}}(\Gamma^{0}(p){\cap}\Gamma(N))\\ \left(\begin{smallmatrix}d&0\\ 0&1\\ \end{smallmatrix}\right)\cdot(f_{a})_{a}=(f_{ad})_{a}\,,\quad g\cdot(f_{a})_{a}=(f_{a}[\tilde{g_{a}}]_{2})_{a}\text{ if }\det g=1\,,\end{gathered}$ (6.5.1)6.5.1( 6.5.1 )
where he operation $[\cdot]_{2}$ is as in ((6.4.1)6.4.1( 6.4.1 )), and $\tilde{g_{a}}$ is chosen as above. Since pullback of differentials along the natural projection $M_{\{\mathrm{Id}\}\times B_{0}(p)}\to M_{H_{p}}$ identifies $H^{0}(\mathcal{A}_{H,p},\Omega^{1})$ with the subspace of $H^{0}(\mathcal{A}_{\{\mathrm{Id}\},p},\Omega^{1})$ made of $H$-invariant differentials, we get the isomorphism
$T_{\ell}\lefttorightarrow H^{0}(\mathcal{A}_{H,p,\mathbb{C}},\Omega^{1})\cong\left(S_{2}^{p\text{-new}}(\Gamma^{0}(p){\cap}\Gamma(N))\otimes_{\mathbb{C}}\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}\right)^{H}\righttoleftarrow\tilde{T}_{\ell}\otimes\sigma_{\ell}$
This, together with Lemma 6.1, Theorem 4.6 and the fact that $A$ is conjugated to $A^{*}$ (Proposition 2.2.2) implies
Theorem 6.5.2
Let $G=G(p,\ell,H)$ be the graph in Definition 1.2, let $\operatorname{Ker}(w_{*}),\operatorname{Ker}(w_{i,*})$ be the subspaces of $\mathbb{C}^{V(G)}$ described in 2.2 and let $S$ be the $p$-new part of $S_{2}(\Gamma^{0}(p){\cap}\Gamma(N))$. Then
[…]
In words $\ker(w_{*})=\oplus_{i}\operatorname{Ker}(w_{i,*})$, as a module over the adjacency matrix of the graph, is isomorphic to the subspace of $S\otimes_{\mathbb{C}}\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}$ fixed by $H$, as a module over $\tilde{T}_{\ell}\otimes\sigma_{\ell}$ (for the action of $H<\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$ see ((6.5.1)6.5.1( 6.5.1 ))).
Remark 6.5.3
In Remark 6.1.6 we pointed out that $M_{H_{p}}$ can be described using either $\Gamma^{0}(p)$ or $\Gamma_{0}(p)$. Following the same lines, Theorem 6.5.2 remains true after substituting $S_{2}^{p\text{-new}}(\Gamma^{0}(p){\cap}\Gamma(N))$ with
$S_{2}^{p\text{-new}}(\Gamma_{0}(p){\cap}\Gamma(N)):=\Big{(}S_{2}(\Gamma(N))+S_{2}(\Gamma(N))[\left(\begin{smallmatrix}p&0\\ 0&1\\ \end{smallmatrix}\right)]_{2}\Big{)}^{\perp}\quad\subset S_{2}(\Gamma_{0}(p){\cap}\Gamma(N))\,.$
and after slightly modifying the action of $\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$ in ((6.5.1)6.5.1( 6.5.1 )), i.e. asking that $\tilde{g_{a}}\in\Gamma_{0}(p)$.
We also rephrase Theorem 6.5.2, for certain choices of $H$, using modular forms for
$\Gamma_{1}(k)=\{m\in\mathrm{SL}_{2}(\mathbb{Z}):m\equiv\left(\begin{smallmatrix}1&*\\ 0&1\\ \end{smallmatrix}\right)\bmod k\}\,,\quad\Gamma_{0}(k)=\{m\in\mathrm{SL}_{2}(\mathbb{Z}):m\equiv\left(\begin{smallmatrix}*&*\\ 0&*\\ \end{smallmatrix}\right)\bmod k\}\,.$
Such modular forms received more attention in the literature, e.g. in the asymptotic estimates in [45] which we later use. We use the decomposition, (see [18, Section 4.3, page 119]),
$S_{2}(\Gamma_{1}(k))=\bigoplus_{\chi\in(\mathbb{Z}/k\mathbb{Z})^{\times,\vee}}S_{2}(\Gamma_{1}(k),\chi)$ (6.5.4)6.5.4( 6.5.4 )
where $\chi$ varies across all characters modulo $k$. In particular, it follows from the definitions that $S_{2}(\Gamma_{0}(p)\cap\Gamma_{1}(N))$ is a subspace of $S_{2}(\Gamma_{1}(Np))$ and precisely the subspace fixed by all the diamond operators (in the sense of [18, Section 5.2]) $\langle d\rangle$ for $d\equiv 1\bmod N$. This implies that
$S_{2}(\Gamma_{0}(p)\cap S_{2}(\Gamma_{1}(N))=\bigoplus_{\chi\in(\mathbb{Z}/N\mathbb{Z})^{\times,\vee}}S_{2}(\Gamma_{1}(pN),\chi)\,.$
where we notice that we are not summing over all characters $\chi$ modulo $Np$, as in ((6.5.4)6.5.4( 6.5.4 )), instead we only look at the characters $\chi\colon(\mathbb{Z}/Np\mathbb{Z})^{\times}\to\mathbb{C}^{\times}$ that factor through the projection $(\mathbb{Z}/Np\mathbb{Z})^{\times}\to(\mathbb{Z}/N\mathbb{Z})^{\times}$. Moreover, if $f$ is a modular form in $S_{2}(\Gamma_{1}(N),\chi)$ for some character $\chi$ modulo $N$, then both $f$ and $f[\left(\begin{smallmatrix}p&0\\ 0&1\\ \end{smallmatrix}\right)]_{1}$ belong to $S_{2}(\Gamma_{1}(Np),\chi)$ by [18, Proposition 5.6.2]. Using this fact we define the spaces of $p$-new forms
$\begin{gathered}S_{2}^{p-\mathrm{new}}(\Gamma_{0}(p)\cap\Gamma_{1}(N)):=\Big{(}S_{2}(\Gamma_{1}(N))+S_{2}(\Gamma_{1}(N))[\left(\begin{smallmatrix}p&0\\ 0&1\\ \end{smallmatrix}\right)]_{2}\Big{)}^{\perp}\quad\subset S_{2}(\Gamma_{0}(p){\cap}\Gamma_{1}(N))\,,\\ S_{2}^{p-\mathrm{new}}(\Gamma_{1}(pN),\chi):=\Big{(}S_{2}(\Gamma_{1}(N),\chi)+S_{2}(\Gamma_{1}(N),\chi)[\left(\begin{smallmatrix}p&0\\ 0&1\\ \end{smallmatrix}\right)]_{2}\Big{)}^{\perp}\quad\subset S_{2}(\Gamma_{1}(pN),\chi)\,,\end{gathered}$
where $\chi$ is modulo $N$ and the orthogonal is taken with respect to the Petersson inner product.
Theorem 6.5.5
Let $G(p,\ell,H)$ be the graph in Definition 1.2, with vertices $V$ and adjacency matrix $A$, and let $\operatorname{Ker}(w_{1,*}),\ldots,\operatorname{Ker}(w_{n,*})$ be the subspaces of $\mathbb{C}^{V}$ described in 2.2.
Then
  • if $H=\{\mathrm{Id}\}<\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$, each $\operatorname{Ker}(w_{i,*})$, as an $A$-module, is isomorphic to $S^{\prime}\otimes_{\mathbb{C}}\mathbb{C}^{L}$, as a module over $\tilde{T}_{\ell}\otimes\sigma_{\ell}$, where $L=\langle\ell\rangle\subset(\mathbb{Z}/N\mathbb{Z})^{\times}$, $\sigma_{\ell}\colon\mathbb{C}^{L}\to\mathbb{C}^{L}$ sends $(a_{x})_{x\in L}$ to $(a_{x\ell})_{x\in L}$, and $S^{\prime}$ is the following space of modular forms
    $S^{\prime}=\bigoplus_{\chi\in(\mathbb{Z}/N\mathbb{Z})^{\times,\vee}}S_{2}^{p\text{-new}}(\Gamma_{1}(pN^{2}),\chi)$
    with $\chi$ varying across the characters that factor through the projection $\mathbb{Z}/pN^{2}\mathbb{Z}\to\mathbb{Z}/N\mathbb{Z}$.
  • if $H=B_{0}(N)=\{\left(\begin{smallmatrix}*&0\\ *&*\\ \end{smallmatrix}\right)\}$ then $n=1$ and $\operatorname{Ker}(w_{1,*})=\{(x_{v})_{v}\in\mathbb{C}^{V}:\sum_{v}x_{v}=0\}$, as a module over $A$ is isomorphic to $S_{2}^{p-\mathrm{new}}(\Gamma_{0}(pN))$ as a module over $\tilde{T}_{\ell}$.
  • if $H=B_{1}(N)=\{\left(\begin{smallmatrix}*&0\\ *&1\\ \end{smallmatrix}\right)\}$ then $n=1$ and $\operatorname{Ker}(w_{1,*})=\{(x_{v})_{v}\in\mathbb{C}^{V}:\sum_{v}x_{v}=0\}$, as a module over $A$ is isomorphic to $S^{\prime}$ as a module over $\tilde{T}_{\ell}$, with
    $S^{\prime}=S_{2}^{p-\mathrm{new}}(\Gamma_{0}(p)\cap\Gamma_{1}(N))=\bigoplus_{\chi\in(\mathbb{Z}/N\mathbb{Z})^{\times,\vee}}S_{2}^{p-\mathrm{new}}(\Gamma_{1}(pN),\chi).$
  • if $H$ is a non-split Cartan of level $N$, then $n{=}1$ and $*\operatorname{Ker}(w_{1,*}$ as an $A$-module, is isomorphic to
    $\bigoplus_{d|N}S_{2}^{\mathrm{new}}(\Gamma_{0}(pd^{2}))$
    as a $\tilde{T}_{\ell}$-module (see [18, Section 5.6] for the definition of $S_{2}^{\mathrm{new}}$).
Proof.
By Lemma 6.1 it is enough to describe the $T_{\ell}$-module $H^{0}(\mathcal{A}_{H,p},\Omega^{1})$.
The cases $H=B_{0}(N)=\{\left(\begin{smallmatrix}*&0\\ *&*\\ \end{smallmatrix}\right)\}$ and $B_{1}(N)=\{\left(\begin{smallmatrix}*&0\\ *&1\\ \end{smallmatrix}\right)\}$ can be treated with the same arguments used for the full level structure in Theorem 6.5.2, even slightly easier: $M_{B_{0}(N)_{p}}(\mathbb{C})$ and $M_{B_{1}(N)_{p}}(\mathbb{C})$ are connected and isomorphic to $\Gamma_{0}(pN)\backslash\overline{\mathbb{H}}$ and $(\Gamma_{0}(p){\cap}\Gamma_{1}(N))\backslash\overline{\mathbb{H}}$, and, since $\left(\begin{smallmatrix}\ell&0\\ 0&1\\ \end{smallmatrix}\right)$ belongs to $H$, the graph is connected and then $T_{\ell}$ acts exactly as $\tilde{T}_{\ell}$.
The full level structure case is a consequence of the Hecke-equivariant isomorphisms
$\mathcal{M}_{B^{\prime}(N^{2})}$ $\longrightarrow\mathcal{M}_{N}\,,\qquad\qquad\qquad(E,(P,Q))\longmapsto(E/\langle nQ\rangle,(nP,Q))$
$\mathcal{M}_{B^{\prime}(N^{2})_{p}}$$\longrightarrow\mathcal{M}_{\{\mathrm{Id}\}\times B_{0}(p)}\,,\qquad\qquad\qquad(E,(P,Q),G)\longmapsto(E/\langle nQ\rangle,(nP,Q),G)$
where $B^{\prime}(N^{2})$ is the subgroup $\{\left(\begin{smallmatrix}1+N*&0\\ *&1+N*\\ \end{smallmatrix}\right)\}$ of $\mathrm{GL}_{2}(\mathbb{Z}/N^{2}\mathbb{Z})$ and where we identify isomorphisms $\phi\colon(\mathbb{Z}/k\mathbb{Z})^{2}\to E[k]$ with basis $(P,Q)$ of the group $E[k]$.
We reduced to $B^{\prime}(N^{2})$ structures. The inclusion $B^{\prime}(N^{2})\supset B_{2}(N^{2}):=\{\left(\begin{smallmatrix}1&0\\ *&1\\ \end{smallmatrix}\right)\}$ induces a map $M_{B_{2}(N^{2})}\to M_{B^{\prime}(N^{2})}$ that identifies $H^{0}(M_{B^{\prime}(N^{2})},\Omega^{1})$ with the $B^{\prime}(N^{2})/M_{B_{2}(N^{2})}$-invariant subspace of $H^{0}(M_{B^{\prime}(N^{2})},\Omega^{1})$. Choosing $\{g_{i}\}=\{\left(\begin{smallmatrix}a&0\\ 0&1\\ \end{smallmatrix}\right):a\in(\mathbb{Z}/N^{2}\mathbb{Z})^{\times}\}$, Equation ((6.1.5)6.1.5( 6.1.5 )) gives
$M_{B_{2}(N^{2})}(\mathbb{C})$ $\cong\bigcup_{a\in(\mathbb{Z}/N^{2}\mathbb{Z})^{\times}}\Gamma_{1}(N^{2})\backslash\overline{\mathbb{H}}\,,\quad$ $\big{(}\mathbb{C}/\mathbb{Z}{+}\mathbb{Z}\tau,(\tfrac{a\tau}{N},\tfrac{1}{N})\big{)}\leftrightarrow(\tau,a)\,,$
$M_{B_{2}(N^{2})_{p}}(\mathbb{C})$ $\cong\bigcup_{a\in(\mathbb{Z}/N^{2}\mathbb{Z})^{\times}}\big{(}\Gamma_{1}(N^{2}){\cap}\Gamma_{0}(p)\big{)}\backslash\overline{\mathbb{H}}\,,\quad$ $\big{(}\mathbb{C}/\mathbb{Z}{+}\mathbb{Z}\tau,(\tfrac{a\tau}{N},\tfrac{1}{N}),\langle\tfrac{1}{p}\rangle\big{)}\leftrightarrow(\tau,a)\,.$
The action of $B^{\prime}(N^{2})/B_{2}(N^{2})$ identifies certain components (two points $(\tau,a)$, $(\tau,a^{\prime})$ are identified iff $a\equiv a^{\prime}\pmod{N}$) and that within the same components identifies a point $(\tau,a)$ with the point $(\widetilde{\langle d\rangle}\tau,a)$ for $d\equiv 1\pmod{N}$ and $\widetilde{\langle d\rangle}$ the diamond operator in [18, Section 5.2]. We deduce the following isomorphism of Hecke-modules
$H^{0}(\mathcal{A}_{\{\mathrm{Id}\},p},\Omega^{1})$ $\cong H^{0}(\mathcal{A}_{B_{2}(N^{2}),p},\Omega^{1})^{B^{\prime}(N^{2})/B_{2}(N^{2})}=\bigoplus_{a\in(\mathbb{Z}/N\mathbb{Z})^{\times}}\bigoplus_{\chi\in(\mathbb{Z}/N\mathbb{Z})^{\times,\vee}}S_{2}^{p-\mathrm{new}}(\Gamma_{1}(pN^{2}),\chi)$
$=\left(\bigoplus_{\chi\in(\mathbb{Z}/N\mathbb{Z})^{\times,\vee}}S_{2}^{p-\mathrm{new}}(\Gamma_{1}(pN^{2}),\chi)\right)\otimes\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}$
on which, by the same arguments used in Theorem 6.5.2, the Hecke operator acts as $\tilde{T}_{\ell}\otimes\sigma_{\ell}$.
For $H$ a non-split Cartan our result follows from the $T_{\ell}$-equivariant isogenies [20, Lemma 3.1 and Theorem 3.8]
$\mathrm{Pic}^{0}(M_{H})\sim\prod_{d|N}J_{0}^{\mathrm{new}}(d^{2}))\,,\quad\mathrm{Pic}^{0}(M_{H_{p}})\sim\prod_{d|N}\left(J_{0}^{\mathrm{new}}(d^{2}))^{2}\times J_{0}^{\mathrm{new}}(pd^{2}))\right)\,,$
where $J_{0}^{\mathrm{new}}(k)$ denotes the new part of the Jacobian of $M_{B_{0}(k)}$. ∎
6.6 Automorphisms of the graphs versus automorphisms of spaces modular forms
We now study how the automorphisms in Definitions 3.10, 3.9 and 3.11 act on a point of $M_{\{\mathrm{Id}\}\times B_{0}(p)}$ (or a quotient $M_{H_{p}}$) under the isomorphism ((6.3.1)6.3.1( 6.3.1 )). Recall that a point $(a,\tau)$ corresponds to the elliptic curve $E_{\tau}=\mathbb{C}/\mathbb{Z}{+}\mathbb{Z}\tau$ together with the subgroup $\langle\tfrac{\tau}{p}\rangle$ and the basis $(\tfrac{a\tau}{N},\tfrac{1}{N})$ of $E[N]$ (such a basis corresponds to the isomorphism $\phi_{\tau}\colon(\mathbb{Z}/N\mathbb{Z})^{2}\to E[N]$ sending the standard basis to it).
The Fricke automorphism $\sigma$ sends the point $(a,\tau)$ to the elliptic curve $\mathbb{C}/\mathbb{Z}{+}\mathbb{Z}\tfrac{\tau}{p}$, with the subgroup $\langle\tfrac{1}{p}\rangle$ and with the basis $(\tfrac{a\tau}{N},\tfrac{1}{N})$ of the $N$-torsion. The multiplication by $\tau^{\prime}=-\tfrac{p}{\tau}$ inside $\mathbb{C}$ induces an isomorphism between this elliptic curve and the elliptic curve $E_{\tau^{\prime}}$, with the subgroup $\langle\tfrac{\tau^{\prime}}{p}\rangle$ and the basis $(-\tfrac{ap}{N},\tfrac{\tau^{\prime}}{N})$, namely the point of $(\tau^{\prime},\left(\begin{smallmatrix}0&1\\ -ap&0\\ \end{smallmatrix}\right))$ under the canonical isomorphism ((6.1.3)6.1.3( 6.1.3 )). If we now apply the action ((6.1.2)6.1.2( 6.1.2 )) of a matrix
$\tilde{m}\in\Gamma^{0}(p)\quad\text{ such that }m\equiv\left(\begin{smallmatrix}0&-1\\ 1&0\\ \end{smallmatrix}\right)\bmod N\,,$
we see that this point is equivalent to the point $(m(\tau^{\prime}),\left(\begin{smallmatrix}ap&0\\ 0&1\\ \end{smallmatrix}\right))$, that is the point $(m\left(\begin{smallmatrix}0&-p\\ 1&0\\ \end{smallmatrix}\right)\tau,ap)$. We deduce that
$\sigma^{*}=[\tilde{m}\left(\begin{smallmatrix}0&-p\\ 1&0\\ \end{smallmatrix}\right)]_{2}\otimes\sigma_{p}\quad\text{in }H^{0}(M_{\{\mathrm{Id}\}\times B^{0}(p),\mathbb{C}},\Omega^{1})\cong S_{2}\left(\Gamma(N){\cap}\Gamma^{0}(p)\right)\otimes_{\mathbb{C}}\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}$
where $\sigma_{p}\lefttorightarrow\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}$ is the shift $(x_{a})\mapsto(x_{ap})$. Inspired by the above discussion we give the following
Definition 6.6.1
The Fricke automorphism on full level modular forms is
$w_{p}\colon S_{2}\left(\Gamma(N){\cap}\Gamma^{0}(p)\right)\longrightarrow S_{2}\left(\Gamma(N){\cap}\Gamma^{0}(p)\right)\,,\quad f\longmapsto f[m_{\sigma}]_{2}$
for $m_{\sigma}=\tilde{m}\left(\begin{smallmatrix}0&-p\\ 1&0\\ \end{smallmatrix}\right)$ and $\tilde{m}\in\Gamma^{0}(p)$ a matrix congruent to $\left(\begin{smallmatrix}0&-1\\ 1&0\\ \end{smallmatrix}\right)$ modulo $N$.
For matricial automorphisms as in Definition 3.9 we have already computed their action in Equation ((6.5.1)6.5.1( 6.5.1 )). In particular, diamond operators $\langle d\rangle$ act as $\widetilde{\langle d\rangle}\otimes\sigma_{d^{2}}$ for $\widetilde{\langle d\rangle}$ as in the next definition (which coincides with the diamond operator in [18, Section 5.2])
Definition 6.6.2
Given $H<\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$, for each $d\in(\mathbb{Z}/N\mathbb{Z})^{\times}$, we have a diamond operator
$\widetilde{\langle d\rangle}\colon S_{2}(\Gamma_{H})\longrightarrow S_{2}(\Gamma_{H})\,,\quad f\longmapsto f[\tilde{m}_{d}]_{2}\,,$
for $\tilde{m}_{d}\in\mathrm{SL}_{2}(\mathbb{Z})$ a matrix congruent to $\left(\begin{smallmatrix}d^{-1}&0\\ 0&d\\ \end{smallmatrix}\right)$ modulo $N$.
Let us now suppose that $N=Mq$ for $M,q$ coprime, $q$ a prime power, and that $H=\tilde{H}\times B_{0}(q)$ as in (3.1). Under the canonical isomorphism ((6.1.3)6.1.3( 6.1.3 )), a point $(\tau,\left(\begin{smallmatrix}a&b\\ c&d\\ \end{smallmatrix}\right))\in M_{H_{p}}(\mathbb{C})$ corresponds to the elliptic curve $E_{\tau}=\mathbb{C}/\mathbb{Z}{+}\mathbb{Z}\tau$ together with the subgroups $\langle\tfrac{\tau}{p}\rangle\subset E_{\tau}[p]$ and $\langle\tfrac{b\tau+d}{q}\rangle\subset E_{\tau}[q]$ and the basis $(\tfrac{a\tau+c}{M},\tfrac{b\tau+d}{M})$ of $E_{\tau}[M]$. The image of a point $(\tau,a)$ under the $q$-th Atkin-Lehner $w_{q}$ is the elliptic curve $\mathbb{C}/\mathbb{Z}\tfrac{1}{q}{+}\mathbb{Z}\tau$ together with the subgroups $\langle\tfrac{\tau}{p}\rangle$ and $\langle\tfrac{\tau}{q}\rangle$ and the basis $(\tfrac{a\tau}{M},\tfrac{1}{M})$ of the $M$-torsion, which, for $\tau^{\prime}=p\tau$ is isomorphic (under the map $z\to qz)$ to the the elliptic curve $\mathbb{C}/\mathbb{Z}{+}\mathbb{Z}\tau^{\prime}$ together with the subgroups $\langle\tfrac{\tau}{{}^{\prime}}p\rangle$ and $\langle\tfrac{\tau}{{}^{\prime}}q\rangle$ and the basis $(\tfrac{a\tau^{\prime}}{M},\tfrac{q}{M})$ of the $M$-torsion. This last datum corresponds to a point $(q\tau,m)$ for $m\in\mathrm{GL}_{2}(\mathbb{Z}/qM\mathbb{Z})$ that is congruent to $\left(\begin{smallmatrix}a&0\\ 0&q\\ \end{smallmatrix}\right)$ modulo $M$ and congruent to (***0)$\ *&0\\ \end{smallmatrix}\right)$ modulo $q$. If we apply the action ((6.1.2)6.1.2( 6.1.2 )) by a matrix
$\tilde{m}_{q}\in\Gamma^{0}(p)\quad\text{such that }\tilde{m}_{q}\equiv\left(\begin{smallmatrix}q&0\\ 0&q^{-1}\\ \end{smallmatrix}\right)\bmod M\,,\tilde{m}_{q}\equiv\left(\begin{smallmatrix}0&-1\\ 1&0\\ \end{smallmatrix}\right)\bmod q\,,$ (6.6.3)6.6.3( 6.6.3 )
the same point is moved to the point $(\tilde{m}_{q}\left(\begin{smallmatrix}q&0\\ 0&1\\ \end{smallmatrix}\right)\tau,\left(\begin{smallmatrix}a(q+M)&0\\ 0&1\\ \end{smallmatrix}\right))$. We deduce that
$w_{q}^{*}=[\tilde{m}_{q}\left(\begin{smallmatrix}q&0\\ 0&1\\ \end{smallmatrix}\right)]_{2}\otimes\sigma_{q+M}\lefttorightarrow\left(S_{2}\left(\Gamma(N){\cap}\Gamma^{0}(p)\right)\otimes_{\mathbb{C}}\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}\right)^{H}$ (6.6.4)6.6.4( 6.6.4 )
where $\sigma_{q+M}\lefttorightarrow\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}$ is the shift $(x_{a})\mapsto(x_{a(q+M)})$.
This discussion, together with Proposition 4.9, Theorem 4.6, and Lemma 6.1 implies the following result. Notice that by Remark 6.2 the automorphisms act by pushforward, or equivalently by pullback of their inverses, on the 1-forms.
Theorem 6.6.5
Let $G=G(p,\ell,H)$ be the graph in Definition 1.2, with $V$ the set of vertices and $\operatorname{Ker}(w_{1,*}),\ldots,\operatorname{Ker}(w_{n,*})$ the subspaces of $\mathbb{C}^{V}$ described in Remark 2.2.
Then there is an isomorphism
$\bigoplus_{i=1}^{n}\operatorname{Ker}(w_{i,*})\cong\left(S_{2}^{p\text{-new}}(\Gamma^{0}(p){\cap}\Gamma(N))\otimes_{\mathbb{C}}\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}\right)^{H}$
that simultaneously intertwines the action of the adjoint of the adjacency matrix $A^{*}$ (see also Proposition 2.2.2), the matricial automorphisms $\langle g\rangle$ in Definition 2.1.1, the Galois action in Definition 2.1.2 and, if there, the Atkin-Lehner involutions $w_{q}$ in Definition 3.11 on the left, with the action of $\tilde{T}_{\ell}\otimes\sigma_{\ell}$, the action of a matrix $g^{-1}$ in ((6.5.1)6.5.1( 6.5.1 )), the map $w_{p}\otimes\sigma_{1/p}$ (see Definition 6.6.1) and, if there, the inverse of the map ((6.6.4)6.6.4( 6.6.4 )) on the right (we denote $\sigma_{d}\lefttorightarrow\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}$ is the shift $(x_{a})_{a}\mapsto(x_{ad})_{a}$).
In some special cases we can be slightly more explicit.
Theorem 6.6.6
Keep the notation as in Theorem 6.5.5 and let $A^{*}$ be the adjoint of the adjacency matrix, as in Proposition 2.2.2.
  • if $H=\{\mathrm{Id}\}$, then $\oplus_{i}\operatorname{Ker}(w_{i,*})$, as module over $A^{*}$, over the Galois action, and over the diamond operators $\langle d\rangle$, is isomorphic to $S^{\prime}\otimes_{\mathbb{C}}\mathbb{C}^{(\mathbb{Z}/N\mathbb{Z})^{\times}}$, as a module over $\tilde{T}_{\ell}\otimes\sigma_{\ell}$, over $w_{p}\otimes\sigma_{1/p}$ and over $\widetilde{\langle d\rangle}^{-1}\otimes\sigma_{d^{-2}}$.
  • if $H=B_{0}(N)=\{\left(\begin{smallmatrix}*&0\\ *&*\\ \end{smallmatrix}\right)\}$ then $n=1$ and $\operatorname{Ker}(w_{1,*})$, as a module over $A^{*}$, over the Galois action, and over the Atkin-Lehner involutions $w_{q}$, is isomorphic to $S_{2}^{p-\mathrm{new}}(\Gamma_{0}(pN))$ as a module over $\tilde{T}_{\ell}$, over the Fricke involution $w_{p}$, and over the other Atkin-Lehner involutions $w_{q}$ in [6].
  • if $H=B_{1}(N)=\{\left(\begin{smallmatrix}*&0\\ *&1\\ \end{smallmatrix}\right)\}$ then $n=1$ and $\operatorname{Ker}(w_{)}{1,*})$, as a module over $A^{*}$, over the Galois action, and over the diamond operators $\langle d\rangle$, is isomorphic to $S^{\prime}$, as a module over $\tilde{T}_{\ell}$, over $w_{p}$ and over $\widetilde{\langle d^{-1}\rangle}$.
  • if $H$ is a non-split Cartan, then $n=1$ and $\operatorname{Ker}(w_{1,*})$ as a module over $A^{*}$, over the Galois action, and over the nontrivial matricial automorphisms $\langle g_{q}\rangle$ for $q^{e}$ a prime power in the factorization of $N$ and $g_{q}$ the only elements in the normalizer of $H$ such that $g_{q}\equiv\mathrm{Id}\pmod{N/q^{e}}$, is isomorphic to $\bigoplus_{d|N}S_{2}^{\mathrm{new}}(\Gamma_{0}(pd^{2}))$ as a module over $\tilde{T}_{\ell}$-module, over the $p$-th Atkin Lehner involution (see [6]) and over the $q$-th Atkin-Lehner involution (that acts trivially on $S_{2}^{\mathrm{new}}(\Gamma_{0}(pd^{2}))$ when $q\nmid d$)
Remark 6.3
To have an isomorphism which is simultaneously equivariant with respect to all automorphisms, in Theorems 6.6.5 and 6.6.6 we used the adjoint $A^{*}$ of the adjacency matrix. Instead, in Theorems 6.5.2 and 6.5.5, for merely aesthetic reasons, we preferred using the adjacency matrix, which is conjugated to its adjoint.
6.7 Asymptotic distribution of the eigenvalues
Following Serre [45], given a linear diagonalizable operator $P$ with spectrum $\sigma(P)$ and domain $V$ of finite dimension $r$, and an we introduce the probability measure
$\mu(P,V):=\frac{1}{r}\sum_{\lambda\in\sigma(P)}\delta_{\lambda}$
where $\delta_{\lambda}$ is a Dirac mass at $\lambda$. Let us also recall the Kensten-McKay measure supported on the Hasse interval $[-2\sqrt{\ell},2\sqrt{\ell}]$ from Equation (1.10)
$\mu_{\ell}=\frac{\ell+1}{\pi}\frac{\sqrt{\ell-x^{2}/4}}{\ell(\ell^{1/2}+\ell^{-1/2})^{2}-x^{2}}dx$
We are interested in $\mu(P,V)$ when $P$ is a Hecke operator and $V$ is one of the spaces appearing in Theorem 6.5.5. The following theorem gives asymptotics, implying Corollary 1.11.
Theorem 6.7.1
Fix a prime $\ell$, a positive integer $N$ coprime with $\ell$, and let $p_{i}$ be an increasing sequence of prime numbers coprime with $N\ell$. Then
$\lim_{i\to\infty}\mu\left(T_{\ell},S_{2}^{p_{i}-new}(\Gamma_{0}(p_{i}N))\right)=\lim_{i\to\infty}\mu\left(T_{\ell},\bigoplus_{d|N}S_{2}^{\mathrm{new}}(\Gamma_{0}(p_{i}d^{2}))\right)=\mu_{\ell}\,,$
and, for each character $\chi$ modulo $N$,
$\lim_{i\to\infty}\mu\left(T_{\ell},S_{2}^{p_{i}\text{-new}}(\Gamma_{1}(p_{i}N),\chi)\right)=\lim_{i\to\infty}\mu\left(T_{\ell},S_{2}^{p_{i}\text{-new}}(\Gamma_{1}(p_{i}N^{2}),\chi)\right)=\sqrt{\chi(\ell)}\mu_{\ell}\,.$
Observe that $\mu_{\ell}=-\mu_{\ell}$, so it does not matter which sign of the square root of $\chi(\ell)$ we choose.
Proof.
Let us first prove the theorem for $S_{2}^{p_{i}-new}(\Gamma_{0}(p_{i}N))$. As Hecke modules we have
$S_{2}(\Gamma_{0}(p_{i}N))=S_{2}^{p_{i}-\mathrm{new}}(\Gamma_{0}(p_{i}N))\oplus S_{2}(\Gamma_{0}(N))^{\oplus 2}.$
Passing to measures, and denoting $d(k)=\dim S_{2}(\Gamma_{0}(k))$, $d(p,k)=\dim S_{2}^{p\text{-new}}(\Gamma_{0}(pk))$, we get
$\mu(T_{\ell},S_{2}(\Gamma_{0}(p_{i}N))=\frac{d(p_{i},N)}{d(p_{i}N)}\mu(T_{\ell},S_{2}(\Gamma_{0}(p_{i}N))^{p_{i}-new})+2\frac{d(N)}{d(p_{i}N)}\mu(T_{\ell},S_{2}(\Gamma_{0}(N)))$
the second addendum on the right hand side goes to zero when $i$ goes to infinity, hence we deduce the claim from [45, Theorem 1].
The other cases are proved in the same way, replacing [45, Theorem 1] first with [45, Theorem 1] and then [45, Theorem 4].
Appendix A Correspondences on nodal curves
In the first part of this Appendix we recall for the reader convenience well-known facts and notations about the Picard group of modular curves. We then use it to state and prove Proposition A.7.
Suppose we are given two smooth projective curves $C_{1},C_{2}$ over a field $k=\overline{k}$. We allow for $C_{1}$ and $C_{2}$ to be disconnected, so let’s keep track of the components $C_{1}^{1},\ldots,C_{1}^{r}$ of $C_{1}$, and $C_{2}^{1},\ldots,C_{2}^{r}$. We suppose that for each $j=1,\ldots,r$ we are given distinct points $x_{1}^{j},\ldots x_{n_{j}}^{j}\in C_{1}^{j}(k)$ and $y_{1}^{j},\ldots y_{n_{j}}^{j}\in C_{2}^{j}(k)$, and we look at the nodal curve
$X=(C_{1}\sqcup C_{2})/x_{i}^{j}=y_{i}^{j}\,,$ (A.1)
We notice that $X$ has $r$ connected components, namely the curves $X_{j}=(C_{1}^{j}\sqcup C_{2}^{j})/x_{i}^{j}=y_{i}^{j}$, each one having $2$ irreducible components.
Let $J=\mathrm{Pic}^{0}_{X/k}$ be the scheme representing invertible sheaves on $X$ having degree 00 when restricted to each irreducible component of $X$. In particular the natural maps $C_{1}\to X$ and $C_{2}\to X$ induce by pull back a map
$J\longrightarrow\mathrm{Pic}^{0}_{C_{1}/k}\times\mathrm{Pic}^{0}_{C_{2}/k}\,.$ (A.2)
Such a map is surjective: given invertible sheaves $\mathcal{L}_{i}$ over $(C_{i})_{l}$, we can construct a (non-canonical) lift of $(\mathcal{L}_{1},\mathcal{L}_{2})$ by choosing generators $v_{i}^{j},w_{i}^{j}$ of $(x_{i}^{j})^{*}\mathcal{L}_{1},(y_{i}^{j})^{*}\mathcal{L}_{2}$ and defining the invertible sheaf $\mathcal{L}=\mathcal{L}_{\mathcal{L}_{1},\mathcal{L}_{2},(v_{i}^{j},w_{i}^{j})_{i,j}}$ on $X$ associating to each open $U\subset X$, the module
$\mathcal{L}(U)=\{(f,g)\in\mathcal{L}_{1}(U\cap C_{1})\times\mathcal{L}_{2}(U\cap C_{2}):\,f(x_{i}^{j})/v_{i}^{j}=g(y_{i}^{j})/w_{i}^{j}\text{ for each }i,j\}\,.$ (A.3)
We notice that the structure sheaf is a particular case of the above construction, namely when $\mathcal{L}_{i}=\mathcal{O}_{C_{i}}$ and $v_{i}=x_{i}^{*}1,w_{i}=y_{i}^{*}1$. Moreover, all the lifts of $(\mathcal{L}_{1},\mathcal{L}_{2})$ are obtained with this construction: given a lift $\mathcal{M}$, we choose for each $i$ a section trivializing $s_{i}$ of $\mathcal{M}_{x_{i}}$, which determines by pull back sections $v_{i},w_{i}$; then the pull back of sections to $C_{i}$ determines a morphisms of $\mathcal{O}$-modules $\mathcal{M}\to\mathcal{L}_{\mathcal{L}_{1},\mathcal{L}_{2},(v_{i},w_{i})_{i}}$, which is an isomorphisms because of how the structure sheaf is defined.
Since map (A.2) is surjective, we have an exact sequence of group schemes over $k$
$0\longrightarrow T\longrightarrow J\longrightarrow\mathrm{Pic}^{0}_{C_{1}/k}\times\mathrm{Pic}^{0}_{C_{2}/k}\longrightarrow 0\,,$ (A.4)
for a certain group scheme $T$. For every $k$-algebra $A$ we can describe the points on $T$ explicitly using (A.3): for every choice of $i,j$, the line bundle $(y_{i}^{j})_{\operatorname{Spec}A}^{*}\mathcal{O}_{C_{2,\operatorname{Spec}A}}$ is canonically trivial, hence hence its generating sections are conanically elements of $A^{\times}$; in particular, every line bundle on $X_{\operatorname{Spec}A}$ that is trivial on the $C_{i}$’s is isomorphic to
$\mathcal{L}_{a}:=\mathcal{L}_{\mathcal{O}_{C_{1}},\mathcal{O}_{C_{2}},(1,a(y_{i}^{j}))}\qquad\text{for some function }a\colon Y=\{y_{1}^{1},\ldots,y_{r}^{n_{r}}\}\longrightarrow A^{\times}\,.$
Which of the invertible sheaves $\mathcal{L}_{a}$ are trivial? Exactly those where $a(y_{i}^{j})$ does not depend on $i$ but only on $j$: indeed $\mathcal{L}_{a}$ is trivial if an only if it is trivial when restricted to each connected component $X^{j}$ of $X$, and, since $\mathcal{L}_{a}|_{X^{j}}$ has degree 00, then it is trivial if and only if it has non trivial global section, which implies that our claim using (A.3) and the fact the only global functions on $C_{1}^{j}$ and $C_{2}^{j}$ are constant. This discussion implies that the following sequence of group schemes over $k$ is exact
[…]
The above exact sequence, allows us to describe the characters of $T$. We have canonical isomorphisms $(\mathbb{G}_{\mathrm{m}}^{Y})^{\vee}=\mathrm{Hom}(\mathbb{G}_{\mathrm{m}}^{Y},\mathbb{G}_{\mathrm{m}})=\mathbb{Z}^{Y}=\bigoplus_{i,j}\mathbb{Z}y_{i}^{j}$ and $(\mathbb{G}_{\mathrm{m}}^{r})^{\vee}=\mathrm{Hom}(\mathbb{G}_{\mathrm{m}}^{r},\mathbb{G}_{\mathrm{m}})=\mathbb{Z}^{r}$ and the map $\Delta$ induces
$\Sigma=\Delta^{\vee}\colon\bigoplus_{i,j}\mathbb{Z}y_{i}^{j}\longrightarrow\mathbb{Z}^{r}\,,\quad\sum_{i,j}m_{i}^{j}y_{i}^{j}\longmapsto\Big{(}\sum_{i=1}^{n_{1}}m_{i}^{1},\ldots,\sum_{i=1}^{n_{r}}m_{i}^{r}\Big{)}\,.$
Then, the exact sequence (A.5) gives the following isomorphism
[…]
In the next proposition we describe how certain correspondences act on $T$ and on its characters, which is applied in the proof of Theorem 4.6 to the Hecke operator 3.6. In the notation of the proposition, we do not keep track of the connected components
Proposition A.7
Let $k$ be an algebraically closed field and let $C=(C_{1}\sqcup C_{2})/(x_{i}=y_{i})_{i=1}^{n}$ and $D=(D_{1}\sqcup D_{2})/(v_{j}=w_{j})_{i=1}^{m}$ be curves over $k$ described as in (A.1), with $C_{i},D_{i}$ smooth.
Let $F,G\colon D\to C$ be maps restricting to $F_{i},G_{i}\colon D_{i}\to C_{i}$ and sending the smooth part of $D$ into the smooth part of $C$ and the nodal points to the nodal points. Then, for each $a\colon\{y_{1},\ldots,y_{n}\}\to k^{\times}$ we have
$G_{*}F^{*}\mathcal{L}_{a}\cong\mathcal{L}_{b}\qquad\text{ for }b:=a\circ{F_{2}}_{*}G_{2}^{*}\colon\,y_{i}\mapsto\prod_{G_{2}(v)=y_{i}}a(F_{2}(v))^{\mathrm{ord}_{v}(G_{2})}\,.$ (A.8)
Where $G_{*}$ is a cycle push-forward.
Let $T$ be the maximal torus of $\mathrm{Pic}_{C/k}$, as in (A.5), and let $T^{\vee}$ be its groups of characters. Keeping track of how the points $y_{i}$ are distributed among the components of $C_{2}$, we get an isomorphism, analogous to (A.6),
$T^{\vee}=\ker\left(\Sigma\colon\bigoplus_{i=1}^{n}\mathbb{Z}y_{i}\to\mathbb{Z}^{r}\right)\,.$
Using the above isomorphism, the map $(G_{*}F^{*})^{\vee}$ is the restriction of the map $H$ below
[…]
Proof.
We first give a description of $T$ in terms of Cartier divisors. For a function $a\colon\{y_{i}\}\to k^{\times}$, take a meromorphic function $f\in k(C_{2})$ such that $f(y_{i})=a(y_{i})$ for every $i$. By (A.3), the pair $(1,f)$ defines a meromorphic section of $\mathcal{L}_{a}$. The divisor associated to this section is supported in $C_{2}\setminus\{v_{1},\dots v_{n}\}$, and can be identified with the divisor $\mathrm{div}f$. As explained for instance in [27, Section 1, Proposition 1.4 (b)], the push-forward of a cycle associated to a meromorphic function can be computed using the norm, so
$G_{*}F^{*}\mathcal{L}_{a}\cong G_{*}F^{*}(\mathrm{div}(1,f))=G_{*}\mathrm{div}(F^{*}(1,f))=\mathrm{div}((1,\mathrm{Norm}_{G_{2}}(F_{2}^{*}f)))=\mathcal{L}_{c}\,,$ (A.10)
for
$c=\mathrm{Norm}_{G_{2}}(F_{2}^{*}f)|_{\{y_{i}\}}\,,$
To prove (A.8), it remains to prove $c=b$. The norm is compatible with pull-backs, i.e. if we want to compute $\mathrm{Norm}_{G_{2}}(F_{2}^{*}f))(y_{i})$ we can look at the base change $G_{2}\colon G_{2}^{-1}(y_{i})\to y_{i}$, the pull-back of $F_{2}^{*}f$ to $G_{2}^{-1}(y_{i})$ and then compute the norm; we conclude that
$(\mathrm{Norm}_{G_{2}}(F_{2}^{*}f))(y_{i})=\prod_{G_{2}(v)=y_{i}}(F_{2}^{*}f)(v)^{\mathrm{ord}_{v}G_{2}}\,.$
Since $G_{2}$ and $F_{2}$ send the smooth part of $D_{2}$ in the smooth part of $C_{2}$ (and analogously for the inverse images), then all the $v$’s appearing above lie in the set $\{w_{j}\}$ and consequently the points $F_{2}(v)$ lie in the set $\{y_{j}\}$, so
$\prod_{G_{2}(v)=y_{i}}(F_{2}^{*}f)(v)^{\mathrm{ord}_{v}G_{2}}=\prod_{G_{2}(v)=y_{i}}f(F_{2}(v))^{\mathrm{ord}_{v}G_{2}}=\prod_{G_{2}(v)=y_{i}}a(F_{2}(v))^{\mathrm{ord}_{v}G_{2}}\,,$
For the second part of the proposition, namely Equation (A.9), it is enough proving that for each $i,j$ we have $(G_{*}F^{*})^{\vee}(y_{i}-y_{j})=H(y_{i}-y_{j})$, which is true since
$(G_{*}F^{*})^{\vee}(y_{i}-y_{j})(\mathcal{L}_{a})=(y_{i}-y_{j})(\mathcal{L}_{b})=\tfrac{b(y_{i})}{b(y_{j})}=\prod_{G_{2}(v)=y_{i}}\!\!a(F_{2}(v))^{\mathrm{ord}_{v}(G_{2})}\cdot\prod_{G_{2}(v)=y_{j}}\!\!a(F_{2}(v))^{-\mathrm{ord}_{v}(G_{2})}$
$\quad=a\left(\sum_{G_{2}(v)=y_{i}}\mathrm{ord}_{v}(G_{2})F_{2}(v)-\sum_{G_{2}(v)=y_{j}}\mathrm{ord}_{v}(G_{2})F_{2}(v)\right)=a\left(H(y_{i}-y_{j})\right)=H(y_{i}-y_{j})(\mathcal{L}_{a})\,.$